### v3

# Darboux’s theorem in $p$-adic symplectic geometry

Abstract.
Let $p$ be a prime number. We derive an analog of Moser’s Path Method for $p$-adic analytic manifolds and use it to prove a $p$-adic analog of Darboux’s Theorem. Using it as a stepping stone we give a classification of second-countable $p$-adic analytic symplectic manifolds in terms of $p$-adic volume. This is a symplectic version of a classical result of Serre in $p$-adic analytic geometry from 1965. We also prove a $p$-adic version of Weinstein’s generalization of Darboux’s Theorem to neighborhoods of compact manifolds. Finally we find explicitly Darboux’s coordinates for the physical models by Ablowitz-Ladik and Salerno of the Discrete Nonlinear Schrödinger equation.
Symplectic geometry is deeply connected with classical and quantum mechanics, and a main motivation of this paper is to pursue a $p$-adic version of symplectic geometry following recent developments which incorporate the $p$-adic numbers into mathematical and theoretical physics.
1. Introduction
In this paper we prove $p$-adic analogs of two fundamental results in geometry: Darboux’s local classification of symplectic manifolds (1882) and Serre’s classification of compact $p$-adic analytic manifolds (1965). In fact we will extend Serre’s classification to the noncompact case as well, only assuming that $p$-adic analytic manifolds are second-countable. Our “symplectic” Serre’s classification implies that all second-countable $p$-adic analytic symplectic manifolds admit global standard Darboux coordinates, which stands in very strong contrast with the real case.
We prove these results using a $p$-adic version of the Moser’s Path Method, which we later also apply to study the physical models by Ablowitz-Ladik and Salerno of the Discrete Nonlinear Schrödinger equation.
In fact, symplectic geometry originated in the study of planetary motions and is deeply connected with classical mechanics. A main motivation for the present paper is to pursue a $p$-adic version of sympletic geometry, in the spirit of other important developments in mathematical physics involing the $p$-adic numbers, see Brekke–Freund [2], Dragovich–Khrennikov–Kozyrev–Volovich [14], Dragovich–Khrennikov–Kozyrev–Volovich–Zelenov [15], Vladimirov–Volovich [42], and the references therein. For treatments involving the $p$-adic numbers in string theory, see Brekke–Freund–Olson–Witten [3], Freund–Olson [17], Freund–Witten [18], Fuquen-Tibatá–García-Compeán–Zúñiga-Galindo [19], García-Compeán–López [20], and Volovich [43]. For applications to the theory of black holes, see Chen–Liu–Hung [7]. For a definition of the $p$-adic symplectic group, Heisenberg group and Maslov index, see Hu–Hu [24] and Zelenov [48].
$1.$ $p$-adic Moser’s Path Method
The following is a $p$-adic analytic analog of Moser’s Path Method; see Appendix B for a review of Moser’s Path Method and an explanation of why it cannot be directly applied in the $p$-adic case (the statement below has additional assumptions).
Theorem A ($p$-adic analytic Moser’s Path Method).
Let $\ell$ and $k$ be integers with $0\leqslant k\leqslant\ell$ and $\ell\geqslant 1$. Let $p$ be a prime number. Let $\mathbb{Z}_{p}$ denote the $p$-adic integers. Let $d=2$ if $p=2$ and otherwise $d=1$. Let $M$ be an $\ell$-dimensional $p$-adic analytic manifold and let $Q$ be a compact submanifold of $M$. Let $a_{S}(x,t)$ be a $p$-adic power series in $t$, for every $x\in M$ and $S\subset\{1,\ldots,\ell\}$ with $|S|=k$. Let $(x_{1},\ldots,x_{\ell})$ be the standard $p$-adic coordinates on $(\mathbb{Q}_{p})^{\ell}$. Let $\{\omega_{t}\}_{t\in p^{-d}\mathbb{Z}_{p}}$ be a family of $p$-adic analytic $k$-forms on $M$ given by
$(\omega_{t})_{x}=\sum_{\begin{subarray}{c}S\subset\{1,\ldots,\ell\}\\ |S|=k\end{subarray}}a_{S}(x,t)\bigwedge_{i\in S}\mathrm{d}x_{i}\quad\forall t\in p^{-d}\mathbb{Z}_{p},\forall x\in M.$
Let $\mathrm{L}$ denote the $p$-adic Lie derivative and suppose that the differential equation
$\frac{\mathrm{d}}{\mathrm{d}t}\omega_{t}+\mathrm{L}_{X_{t}}\omega_{t}=0$
has a solution $X_{t}\in\mathfrak{X}(M)$ for every $t\in p^{-d}\mathbb{Z}_{p}$ which satisfies the following conditions:
  1. (i)
    the components of $X_{t}$ are given by $p$-adic power series converging in $M$, in some local coordinates;
  2. (ii)
    for all $m\in Q$, when the $p$-adic power series in (i) are translated so that the origin is $m$, they do not contain terms with degree less than $2$, that is, $X_{t}(m)=0$ and the partial derivatives of the components of $X_{t}$ at $m$ are 0, for all $m\in Q$ and $t\in p^{-d}\mathbb{Z}_{p}$.
Then there exists an open subset $U\subset M$ and a $p$-adic analytic family of $p$-adic analytic diffeomorphisms $\{\psi_{t}:U\to\psi_{t}(U)\subset M\}_{t\in\mathbb{Z}_{p}}$, such that $\psi_{t}^{*}\omega_{t}=\omega_{0}|_{U}$.
1.2. Applications of $p$-adic Moser’s Path Method to symplectic geometry
1.2.1. Applications to local symplectic geometry
Theorem A can be used, as happens in the real case, to prove Darboux’s Theorem in the $p$-adic context:
Theorem B ($p$-adic analytic Darboux’s Theorem).
Let $n$ be a positive integer. Let $p$ be a prime number. Let $(M,\omega)$ be a $2n$-dimensional $p$-adic analytic symplectic manifold and $m\in M$. There exist local $p$-adic analytic coordinates $(x_{1},y_{1},\ldots,x_{n},y_{n})$ around $m$ such that $\omega=\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}.$
Figure 1. Illustration of the $p$-adic analytic Darboux’s Theorem (Theorem B).
It follows from Theorem B that the only local invariant of $p$-adic analytic symplectic manifolds is the dimension. See Figure 1 for an illustration of Theorem B. Using the $p$-adic Moser’s Path Method we can generalize Theorem B to neighborhoods of compact submanifolds, and obtain a $p$-adic analog of the Darboux-Weinstein’s Theorem [45, Theorem 4.1].
Theorem C ($p$-adic analytic Darboux-Weinstein’s Theorem).
Let $n$ be a positive integer. Let $p$ be a prime number. Let $M$ be a $2n$-dimensional $p$-adic analytic manifold and let $Q$ be a compact submanifold of $M$. Let $\omega_{0}$ and $\omega_{1}$ be two $p$-adic analytic symplectic forms on $M$ which coincide on $Q$. Then there exist open neighborhoods $U_{0},U_{1}$ of $Q$ and a $p$-adic analytic diffeomorphism $\psi:U_{0}\to U_{1}$ which fixes $Q$ and such that $\psi^{*}\omega_{1}=\omega_{0}$.
1.2.2. Applications to global symplectic geometry
In what follows, when we say that two $p$-adic manifolds are symplectomorphic, it is assumed that they are so by means of a $p$-adic analytic symplectomorphism.
Theorem D.
Let $v$ be a positive real number. Let $d$ be a positive integer. Then the following statements hold:
  1. (1)
    There exists a noncompact $d$-dimensional $p$-adic analytic manifold $\mathcal{M}_{d,v}\subset(\mathbb{Q}_{p})^{d}$ with $p$-adic volume $v$, defined as follows: if $v=\sum_{k=-\infty}^{m}a_{k}p^{dk}$ is the expression of $v$ in base $p^{d}$ with infinitely many nonzero digits (for example $1$ in base $9$ becomes $0.888\ldots$), then,
    $\mathcal{M}_{d,v}=\bigcup_{k=-\infty}^{m}\bigcup_{i=1}^{a_{k}}\mathrm{B}_{ki}\subset(\mathbb{Q}_{p})^{d}$
    where $\mathrm{B}_{ki}$ is the $d$-dimensional $p$-adic ball defined by
    $\mathrm{B}_{ki}=\mathrm{B}((i_{1}p^{-k-1},i_{2}p^{-k-1},\ldots,i_{d}p^{-k-1}),p^{k}),$
    which are all pairwise disjoint, with $i=\sum_{j=0}^{d-1}i_{j+1}p^{j}$ and $0\leqslant i_{j}\leqslant p-1$. See Figure 5 for the balls $\mathrm{B}_{ki}$.
  2. (2)
    If, in addition, $v$ is rational and its denominator is a power of $p$, then there also exists a compact $d$-dimensional $p$-adic analytic manifold $\mathcal{M}_{d,v}^{\prime}\subset(\mathbb{Q}_{p})^{d}$ with $p$-adic volume $v$, defined as follows: if $v=\sum_{k=\ell}^{m}a_{k}p^{dk}$ is the expression of $v$ in base $p^{d}$ with finitely many nonzero digits, which exists because $v$ is rational and its denominator is a power of $p$, then
    $\mathcal{M}_{d,v}^{\prime}=\bigcup_{k=\ell}^{m}\bigcup_{i=1}^{a_{k}}\mathrm{B}_{ki}\subset(\mathbb{Q}_{p})^{d}.$
Endow $\mathcal{M}_{d,v}$ and $\mathcal{M}_{d,v}^{\prime}$ with the standard symplectic form $\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}\xi_{i}$. If in the definition of $\mathcal{M}_{d,v}$ and $\mathcal{M}_{d,v}^{\prime}$ we change the balls by other pairwise disjoint balls of the same radius the resulting $p$-adic analytic manifolds are symplectomorphic.
Now we turn to the symplectic version of Serre’s result Theorem C.1. The $p$-adic volume which we use here is the $p$-adic measure with respect to the form $\omega^{n}$, where $\omega$ is the symplectic form, analogously to the definition in the real case.
Theorem E (Symplectic Serre’s classification of $p$-adic analytic manifolds).
Let $n$ be a positive integer. Let $p$ be a prime number. Let $(M,\omega)$ be a second-countable $2n$-dimensional $p$-adic analytic symplectic manifold. Let $v$ be the volume of $(M,\omega)$ and let $\mathcal{M}_{2n,v}$ and $\mathcal{M}_{2n,v}^{\prime}$ be as defined in Theorem D. Endow $(\mathbb{Q}_{p})^{2n}$, $\mathcal{M}_{2n,v}$ and $\mathcal{M}_{2n,v}^{\prime}$ with the standard symplectic form $\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}\xi_{i}$.
  1. (1a)
    If $M$ is noncompact and $v=\infty$, $(M,\omega)$ is symplectomorphic to $(\mathbb{Q}_{p})^{2n}$.
  2. (1b)
    If $M$ is noncompact and $v$ is finite, $(M,\omega)$ is symplectomorphic to $\mathcal{M}_{2n,v}$.
  3. (1c)
    If $M$ is compact, then $v$ is a rational number with a power of $p$ as denominator and $(M,\omega)$ is symplectomorphic to $\mathcal{M}_{2n,v}^{\prime}$.
  4. (2)
    $\mathcal{M}_{2n,v}$ and $\mathcal{M}_{2n,v^{\prime}}$ are symplectomorphic if and only if $v=v^{\prime}$. The same happens with $\mathcal{M}_{2n,v}^{\prime}$ and $\mathcal{M}_{2n,v^{\prime}}^{\prime}$.
Recall that $\mathcal{M}_{2n,v}$ and $\mathcal{M}_{2n,v}^{\prime}$ are defined as unions of balls, hence Theorem E is a symplectic version of Serre’s Classification (Theorem C.1); the difference is that the symplectic case needs an unbounded number of balls. Theorem E implies that there exists a $p$-adic analytic symplectic embedding $\varphi:(M,\omega)\to((\mathbb{Q}_{p})^{2n},\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}\xi_{i}).$ Hence, there always exist global Darboux coordinates.
Theorem F.
Let $n$ be a positive integer. Let $p$ be a prime number. Let $(M,\omega)$ and $(M^{\prime},\omega^{\prime})$ be second-countable $2n$-dimensional $p$-adic analytic symplectic manifolds with the same $p$-adic volume, finite or infinite, and which are both compact or both noncompact. Then $(M,\omega)$ is symplectomorphic to $(M^{\prime},\omega^{\prime})$.
Theorem F stands in very strong contrast with real symplectic geometry where there is an extensive theory of other invariants, such as symplectic capacities.
1.2.3. Application of the $p$-adic Moser’s Path Method to the Ablowitz-Ladik and Salerno models
In Section 7 we apply the method of proof of Theorem B to a $p$-adic analog of the Ablowitz-Ladik’s Model and Salerno’s Model, which are generalizations of the Discrete Nonlinear Schrödinger Model. In the real case, after changing coordinates so that the symplectic form takes the standard form, the Hamiltonian of these models takes the same form as the mentioned Discrete Nonlinear Schrödinger model (see [5] for details). It is not possible to derive a closed form for the coordinate change, but some intermediate results can be obtained. We will obtain those results for the $p$-adic case (Proposition 7.1), with the same expressions.
Ablowitz-Ladik’s and Salerno’s models use the same nonstandard symplectic form:
$\omega=\mathrm{i}\sum_{j\in J}\frac{\mathrm{d}\psi_{j}\wedge\mathrm{d}\bar{\psi}_{j}}{1+\mu|\psi_{j}|^{2}}.$
In the $p$-adic case, we are able to recover the result from [5], which tells us that this symplectic form can be taken into the standard one $\omega_{\mathrm{std}}=\mathrm{i}\sum_{j\in J}^{n}\mathrm{d}\psi_{j}\wedge\mathrm{d}\bar{\psi}_{j},$ using as a transformation the flow of the vector field
$X_{t}=\frac{1+\nu(x_{1}^{2}+y_{1}^{2})}{1+(1-t)\nu(x_{1}^{2}+y_{1}^{2})}\left(1-\frac{\log(1+\nu(x_{1}^{2}+y_{1}^{2}))}{2\nu(x_{1}^{2}+y_{1}^{2})}\right)\begin{pmatrix}x_{1}\\ y_{1}\end{pmatrix}$.
Structure of the paper
Section 2 proves some intermediate results which we need in order to derive Moser’s Path Method and prove Darboux’s Theorem. Section 3 derives the $p$-adic Moser’s Path Method (Theorem A). Section 4 proves the $p$-adic Darboux’s Theorem (Theorem B). Section 5 proves the $p$-adic Darboux-Weinstein’s Theorem (Theorem C). Section 6 proves Theorems D and E. Finally, Section 7 applies Darboux’s Theorem to the $p$-adic version of the Ablowitz-Ladik and Salerno models. The appendices recall the concepts we need about $p$-adic numbers, $p$-adic analytic functions, $p$-adic analytic manifolds and $p$-adic symplectic geometry (Appendix A), the real Moser’s Path Method (Appendix B) and the classical Serre’s classification (Appendix C).
Funding
The first author is funded by grant PID2022-137283NB-C21 of MCIN/AEI/ 10.13039/501100011033 / FEDER, UE. The second author is funded by a FBBVA (Bank Bilbao Vizcaya Argentaria Foundation) Grant for Scientific Research Projects with title From Integrability to Randomness in Symplectic and Quantum Geometry.
Acknowledgments
The second author thanks the Dean of the School of Mathematical Sciences Antonio Brú and the Chair of the Department of Algebra, Geometry and Topology at the Complutense University of Madrid, Rutwig Campoamor, for their support and excellent resources he is being provided with to carry out the FBBVA project. The second author also thanks Luis Crespo and Francisco Santos for the hospitality during July and August 2025 at the University of Cantabria when part of this paper was written.
2. Technical results on $p$-adic initial value problems, power series and forms on manifolds
In this section we prove four technical lemmas which we will need to prove the main theorems (Theorems A, B, C, D, E, F). We refer to Appendix A for a quick review of the concepts and results from $p$-adic analysis and $p$-adic geometry we need in this section (in particular see Figures 7 and 8 for representations of $p$-adic balls like the ones which appear in Lemma 2.1 below).
Lemma 2.1 (Multivariable $p$-adic initial value problem).
Let $\ell$ be a positive integer. Let $p$ be a prime number. Let $d=2$ if $p=2$ and otherwise $d=1$. Let $x_{0}\in\mathbb{Q}_{p}$, $R$ a $p$-adic absolute value, $U_{1}=\mathrm{B}(x_{0},R)$ and $U_{2}=\mathrm{B}(x_{0},p^{d}R)$. Let $v\in(\mathbb{Q}_{p})^{\ell}$ and let $V$ be a $p$-adic ball in $(\mathbb{Q}_{p})^{\ell}$ centered at $v$. For every $i\in\{1,\ldots,\ell\}$ let $f_{i}\in\mathbb{Q}_{p}[[x,y_{1},\ldots,y_{\ell}]]$ be a $p$-adic power series converging in $U_{2}\times V$ such that:
  • for every $i\in\{1,\ldots,\ell\}$ and every $x\in U_{2}$ we have that
    $f_{i}(x,v)=0$
    and,
  • for all $i,j\in\{1,\ldots,\ell\}$ and every $x\in U_{2}$,
    $\frac{\partial f_{i}}{\partial y_{j}}(x,v)=0.$
Then there exists a $p$-adic ball $V^{\prime}\subset V$ centered at $v$ such that, for any $(y_{01},\ldots,y_{0\ell})\in V^{\prime}$, the $p$-adic initial value problem of the form
(2.1) $\left\{\begin{aligned} &\frac{\mathrm{d}y_{1}}{\mathrm{d}x}=f_{1}(x,y_{1},\ldots,y_{\ell});\\ &\ldots\\ &\frac{\mathrm{d}y_{\ell}}{\mathrm{d}x}=f_{\ell}(x,y_{1},\ldots,y_{\ell});\\ &y_{1}(x_{0})=y_{01},\,\,\ldots,\,\,y_{\ell}(x_{0})=y_{0\ell},\end{aligned}\right$.
has a unique solution given by $p$-adic power series $y_{1},\ldots,y_{\ell}\in\mathbb{Q}_{p}[[x]]$ which converge in $U_{1}$.
Proof.
Step 1: reduction to a canonical form. By translating the power series if necessary, we may assume that $x_{0}=0$ and $v=0$. This does not change the convergence radii of the series, because the convergence radius of a $p$-adic power series is invariant by translation. After translating, $f_{i}(x,0,\ldots,0)$ is 0 for every $i\in\{1,\ldots,\ell\}$ and every $x\in U_{2}$, and the same happens with its partial derivatives respect to $y_{j}$, hence the series $f_{i}$ have no terms with degree 0 or $1$ in $y_{1},\ldots,y_{\ell}$.
Now we multiply the variables by some constants so that the balls $U_{2}$ and $V$ become $\mathbb{Z}_{p}$ and $(\mathbb{Z}_{p})^{\ell}$, in particular, $R=p^{-d}$.
Step 2: choice of $V^{\prime}$. Let $c_{i}(s_{0},\ldots,s_{\ell})$ be the coefficient of $f_{i}$ with degree $s_{0},s_{1},\ldots,s_{\ell}$ in $x,y_{1},\ldots,y_{\ell}$. Since $f_{i}$ converges in $(\mathbb{Z}_{p})^{\ell+1}$,
(2.2) $\lim_{s\to\infty}\max\Big\{|c_{i}(s_{0},\ldots,s_{\ell})|_{p}:i\in\{1,\ldots,\ell\},\,\,s_{0},s_{1},\ldots,s_{\ell}\in\mathbb{N},\,\,\sum_{h=0}^{\ell}s_{h}=s\Big\}=0.$
In particular, the set
$\Big\{|c_{i}(s_{0},\ldots,s_{\ell})|_{p}:i\in\{1,\ldots,\ell\},\,\,s_{0},s_{1},\ldots,s_{\ell}\in\mathbb{N}\Big\}$
is bounded. This allows us to define, for each $s\in\mathbb{N}$,
$C_{s}=\max\Big\{|c_{i}(s_{0},\ldots,s_{\ell})|_{p}:i\in\{1,\ldots,\ell\},\,\,s_{0},s_{1},\ldots,s_{\ell}\in\mathbb{N},\,\,\sum_{h=1}^{\ell}s_{h}=s\Big\}.$
Since $f_{i}$ has no constant or linear terms in the variables $y_{j}$, we have that
$C_{0}=C_{1}=0.$
Also, (2.2) implies that
(2.3) $\lim_{s\to\infty}C_{s}=0.$
Expression (2.3) implies that only a finite number of the $C_{s},s\in\mathbb{N},$ can be greater than $1$. Hence we may take $r$ small enough so that
$r^{s-1}C_{s}\leqslant 1$
for all $s\geqslant 2$. We define $V^{\prime}=\mathrm{B}(0,r).$
Step 3: uniqueness of solution. We now check that the solution to (2.1) is indeed unique for an initial value in $V^{\prime}$. Let $(y_{01},\ldots,y_{0\ell})\in V^{\prime}$. Since the solution $(y_{1},\ldots,y_{\ell})$ must be given by a power series in $x$, we have that
$y_{i}(x)=\sum_{j=0}^{\infty}a_{ij}x^{j},$
for some $a_{ij}\in\mathbb{Q}_{p}$. By (2.1), $a_{i0}=y_{i}(0)=y_{0i}$. Also by (2.1),
$\sum_{j=0}^{\infty}a_{ij}jx^{j-1}=f_{i}\left(x,\sum_{j=0}^{\infty}a_{1j}x^{j},\ldots,\sum_{j=0}^{\infty}a_{nj}x^{j}\right).$
We expand the right-hand side in terms of the coefficients $c_{i}(s_{0},\ldots,s_{\ell})$:
$\sum_{j=0}^{\infty}a_{ij}jx^{j-1}$ $=\sum_{s_{0},s_{1},\ldots,s_{\ell}\in\mathbb{N}}c_{i}(s_{0},\ldots,s_{\ell})x^{s_{0}}\prod_{h=1}^{\ell}\left(\sum_{j=0}^{\infty}a_{hj}x^{j}\right)^{s_{h}}$
$=\sum_{s_{0},s_{1},\ldots,s_{\ell}\in\mathbb{N}}c_{i}(s_{0},\ldots,s_{\ell})x^{s_{0}}\prod_{h=1}^{\ell}\sum_{j_{1},\ldots,j_{s_{h}}\in\mathbb{N}}a_{hj_{1}}\ldots a_{hj_{s_{h}}}x^{j_{1}+\ldots+j_{s_{h}}}$
$=\sum_{s_{0},s_{1},\ldots,s_{\ell}\in\mathbb{N}}c_{i}(s_{0},\ldots,s_{\ell})x^{s_{0}}\sum_{\begin{subarray}{c}j_{h1},\ldots,j_{hs_{h}}\in\mathbb{N}\\ h\in\{1,\ldots,\ell\}\end{subarray}}\prod_{h=1}^{\ell}a_{hj_{h1}}\ldots a_{hj_{hs_{h}}}x^{j_{h1}+\ldots+j_{hs_{h}}}$
(2.4) $=\sum_{s=0}^{\infty}\sum_{\begin{subarray}{c}s_{0},s_{1},\ldots,s_{\ell}\in\mathbb{N}\\ s_{1}+\ldots+s_{\ell}=s\end{subarray}}c_{i}(s_{0},\ldots,s_{\ell})x^{s_{0}}\sum_{\begin{subarray}{c}h_{1},\ldots,h_{s}\in\{1,\ldots,\ell\}\\ j_{1},\ldots,j_{s}\in\mathbb{N}\end{subarray}}a_{h_{1}j_{1}}\ldots a_{h_{s}j_{s}}x^{j_{1}+\ldots+j_{s}}.$
Equating the degree $k$ parts at both sides of (2.4) results in
(2.5) $(k+1)a_{i,k+1}=\sum_{s=0}^{\infty}\sum_{\begin{subarray}{c}s_{0},s_{1},\ldots,s_{\ell}\in\mathbb{N}\\ s_{1}+\ldots+s_{\ell}=s\end{subarray}}c_{i}(s_{0},\ldots,s_{\ell})\sum_{\begin{subarray}{c}h_{1},\ldots,h_{s}\in\{1,\ldots,\ell\}\\ j_{1},\ldots,j_{s}\in\mathbb{N}\\ s_{0}+j_{1}+\ldots+j_{s}=k\end{subarray}}a_{h_{1}j_{1}}\ldots a_{h_{s}j_{s}}.$
Hence $a_{i,k+1}$ is given as a polynomial in the coefficients of smaller degree, and the solution is unique.
Step 4: convergence of the solution. It is left to prove that $y_{i}(x)$ converges in $U_{1}$. In order to do this, it is enough to check that
(2.6) $|a_{ij}|_{p}\leqslant\frac{r}{|j!|_{p}}.$
We show this by induction on $j$. For $j=0$, it follows from
$|a_{i0}|_{p}=|y_{0i}|_{p}\leqslant r.$
Now we suppose it holds for $0\leqslant j\leqslant k$ and prove it for $j=k+1$. By equation (2.5), $(k+1)a_{i,k+1}$ is equal to a polynomial in the previous coefficients. The $p$-adic absolute value of this polynomial is at most
$|(k+1)a_{i,k+1}|_{p}\leqslant|c_{i}(s_{0},\ldots,s_{\ell})a_{h_{1}j_{1}}\ldots a_{h_{s}j_{s}}|_{p},$
where
$s=\sum_{h=1}^{\ell}s_{h}$
and
$\sum_{b=1}^{s}j_{b}\leqslant k.$
By the definition of $C_{s}$, the induction hypothesis, and our choice of $r$,
$|(k+1)a_{i,k+1}|_{p}$ $\leqslant C_{s}\frac{r}{|j_{1}!|_{p}}\ldots\frac{r}{|j_{s}!|_{p}}$
$\leqslant\frac{r}{|j_{1}!\ldots j_{s}!|_{p}}.$
It is known that $j_{1}!\ldots j_{s}!$ divides $(j_{1}+\ldots+j_{s})!$, which in turn divides $k!$, and we have that
$|(k+1)a_{i,k+1}|_{p}\leqslant\frac{r}{|k!|_{p}},$
which implies
$|a_{i,k+1}|_{p}\leqslant\frac{r}{|(k+1)!|_{p}}.$
The induction step is complete.
Now equation (2.6) implies that $y_{i}$ converges at least in the same radius as the series with coefficient of degree $j$ equal to $1/j!$. But this is precisely the exponential series, whose convergence radius is $p^{-d}$ (see Appendix A), and the ball with radius $p^{-d}$ is exactly $U_{1}$. ∎
Remark 2.2.
Lemma 2.1 is similar in spirit to our previous result [8, Proposition A.9], but that result was weaker and in particular it only concerned one variable.
Example 2.3.
Consider the differential equation
$y^{\prime}=\frac{y^{2}}{1-x}$
with $x_{0}=0$ and $y(0)=y_{0}$. Its power series converges for $x\in p\mathbb{Z}_{p}$, hence we can take $U_{1}=p\mathbb{Z}_{p}$. Its solution can be computed explicitly:
$y=\frac{1}{\frac{1}{y_{0}}+\log(1-x)}.$
If we solve the equation using the method in the proof of Lemma 2.1, we set $y$ to a power series
$y=\sum_{j=0}^{\infty}a_{j}x^{j}$
and substitute in the equation, obtaining
$\sum_{j=0}^{\infty}a_{j}jx^{j-1}$ $=\frac{1}{1-x}\left(\sum_{j=0}^{\infty}a_{j}x^{j}\right)^{2}$
$=\left(\sum_{h=0}^{\infty}x^{h}\right)\left(\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}a_{i}a_{j}x^{i+j}\right)$
$=\sum_{k=0}^{\infty}x^{k}\sum_{i=0}^{k}\sum_{j=0}^{k-i}a_{i}a_{j}$
which gives us a recurrence relation for the $a_{j}$:
$a_{k+1}=\frac{1}{k+1}\sum_{i=0}^{k}\sum_{j=0}^{k-i}a_{i}a_{j},$
with initial term $a_{0}=y_{0}$. Its first terms are
$\left\{\begin{aligned} a_{0}&=y_{0};\\ a_{1}&=y_{0}^{2};\\ a_{2}&=\frac{y_{0}^{2}}{2}+y_{0}^{3};\\ a_{3}&=\frac{y_{0}^{2}}{3}+y_{0}^{3}+y_{0}^{4}.\end{aligned}\right$.
Lemma 2.1 tells us that this power series converges at least in $U_{2}=p^{1+d}\mathbb{Z}_{p}$.
Lemma 2.4.
Let $p$ be a prime number. Let $r$ be a $p$-adic absolute value. Let $\overline{\mathbb{Q}_{p}}$ be the algebraic closure of $\mathbb{Q}_{p}$. Let $f\in\mathbb{Q}_{p}(t)$ be a rational function with $p$-adic coefficients, and suppose that all zeros of the denominator of $f$ in $\overline{\mathbb{Q}_{p}}$ have $p$-adic norm greater than $r$. Then the $p$-adic power series of $f$ converges in $\mathrm{B}(0,r)$.
Proof.
It is enough to prove it for the case when the numerator is $1$, because multiplying a power series converging in $\mathrm{B}(0,r)$ times a polynomial results in another power series converging in the same ball.
For every $i\in\mathbb{N}$, let $a_{i}$ be the coefficient of degree $i$ in the power series of $f$. Let $d$ the degree of the denominator. For $i\in\{0,\ldots,d\}$, let $b_{i}$ the coefficient of degree $i$ in the denominator of $f$. Let $R$ be the absolute value of the root of the denominator with smaller absolute value. We have that $R>r$. We can derive a recurrence for the $a_{i}$ in terms of the $b_{i}$:
$\left\{\begin{aligned} a_{0}b_{0}&=1;\\ \sum_{i=0}^{d}a_{k-i}b_{i}&=0\text{ for }k\in\mathbb{N}.\end{aligned}\right$.
This implies that
$\left\{\begin{aligned} a_{0}&=\frac{1}{b_{0}};\\ a_{k}&=-\sum_{i=1}^{d}a_{k-i}\frac{b_{i}}{b_{0}}\text{ for }k\in\mathbb{N}.\end{aligned}\right$.
For each $i\in\{0,\ldots,n\}$, the ratio $b_{i}/b_{0}$ can be written as the symmetric polynomial of degree $i$ in the roots of
$\sum_{j=0}^{d}b_{d-j}x^{j},$
which are the inverses of the roots of the denominator, and have absolute value at most $1/R$. This implies that the symmetric polynomial of degree $i$ in those roots has absolute value at most $1/R^{i}$ and
$|a_{k}|_{p}$ $\leqslant\max\left\{\left|a_{k-i}\frac{b_{i}}{b_{0}}\right|_{p}:i\in\{1,\ldots,d\}\right\}$
$\leqslant\max\left\{\frac{|a_{k-i}|_{p}}{R^{i}}:i\in\{1,\ldots,d\}\right\}.$
This implies, by induction on $k$, that
$|a_{k}|_{p}\leqslant\frac{|a_{0}|_{p}}{R^{k}},$
for any $k\in\mathbb{N}$, so
$r^{k}|a_{k}|_{p}\leqslant|a_{0}|_{p}\frac{r^{k}}{R^{k}},$
which tends to 0 when $k$ tends to infinity. Hence the power series of $f$ converges in $\mathrm{B}(0,r)$, as we wanted. ∎
Example 2.5.
Consider the function
$f(x)=\frac{1}{x^{2}+1}.$
If 1-1 is not a square in $\mathbb{Q}_{p}$, $f(x)$ is defined in all $\mathbb{Q}_{p}$. However, its power series
$f(x)=1-x^{2}+x^{4}-x^{6}+\ldots$
converges only in the $p$-adic balls with radius smaller than $1$. Lemma 2.4 explains the situation: the denominator has a root of $p$-adic absolute value $1$ in $\overline{\mathbb{Q}_{p}}$, namely the two roots of 1-1. The same example also works with $\mathbb{R}$ instead of $\mathbb{Q}_{p}$ and $\mathbb{C}$ instead of $\overline{\mathbb{Q}_{p}}$.
The following lemma is a $p$-adic version of the Tubular Neighborhood Theorem [4, page 66].
Definition 2.6 ($p$-adic Tubular Neighborhood).
Let $\ell$ be a positive integer. Let $p$ be a prime number. Let $M$ be an $\ell$-dimensional $p$-adic analytic manifold and $Q$ a compact submanifold of $M$. Let $\mathrm{N}(Q)$ be the normal bundle of $Q$:
$\mathrm{N}(Q)=\Big\{(x,v)\in\mathrm{T}M:\,\,v^{\mathrm{T}}u=0\text{ for all }u\in\mathrm{T}_{x}Q\Big\}.$
For $k\in\mathbb{N}\cup\{0\}$, let
$\mathrm{N}(Q)_{k}=\Big\{(x,v)\in\mathrm{N}(Q):\|v\|_{p}\leqslant p^{-k}\Big\}.$
(The expression $v^{\mathrm{T}}u$ refers to the expression of $u$ and $v$ in local coordinates.) A $p$-adic tubular neighborhood of $Q$ is a neighborhood $U$ of $Q$ in $M$ such that $U$ is diffeomorphic to $\mathrm{N}(Q)_{k}$ by means of a $p$-adic analytic diffeomorphism for some $k\in\mathbb{N}$. See Figure 2 for an example.
Figure 2. $p$-adic tubular neighborhoods of dimension 0, 11 and 22 submanifolds of a dimension 33 manifold as in Definition 2.6.
Lemma 2.7 ($p$-adic analytic Tubular Neighborhood Theorem).
Let $\ell$ be a positive integer. Let $p$ be a prime number. Let $M$ be an $\ell$-dimensional $p$-adic analytic manifold and let $Q$ be a compact $p$-adic analytic submanifold of $M$. Then there exists a $p$-adic analytic tubular neighborhood of $Q$ in $M$.
Proof.
Since every $p$-adic analytic manifold can be written as a disjoint union of domains of charts in the atlas by [8, Corollary 3.2], all of which are $p$-adic balls, the problem reduces to the case when $M$ is a $p$-adic ball. We may assume that this ball is $(\mathbb{Z}_{p})^{\ell}$.
We consider the map $f:\mathrm{N}(Q)_{0}\to M$ given by
$f(x,v)=x+v.$
Its differential sends $(u,v)$ to $u+v$, where $u$ is tangent to $Q$ and $v$ is normal to $Q$, hence it is bijective and $f$ is a local diffeomorphism. In order to finish the proof, we will show that $f|_{\mathrm{N}(Q)_{k}}$ is injective for $k$ big enough, which implies that $f(\mathrm{N}(Q)_{k})$ is diffeomorphic to $\mathrm{N}(Q)_{k}$, and it contains $Q$ because $f(x,0)=x$ for $x\in Q$.
Suppose, for a contradiction, that $f|_{\mathrm{N}(Q)_{k}}$ is not injective for any $k$. Then, for each $k\in\mathbb{N}$, there exist $a_{k}$ and $b_{k}$ in $N_{k}$ such that $a_{k}\neq b_{k}$ and $f(a_{k})=f(b_{k})$. Since $Q$ is compact, $N_{0}$ is also compact, hence we can extract from the sequence $(a_{k},b_{k})_{k\in\mathbb{N}}$ a convergent subsequence. Let $(a,b)$ be its limit.
Now $a$ and $b$ are in the limit of the $N_{k}$, hence they are $(x,0)$ and $(y,0)$ for some $x,y\in Q$. Since $f$ is continuous and $f(a_{k})=f(b_{k})$,
$f(x,0)=f(y,0).$
Together with $f(x,0)=x$ for $x\in Q$, we must have $x=y$.
We know that $f$ is a local diffeomorphism, hence there is a neighborhood $V$ of $(x,0)$ in $N$ such that $f$ is injective when restricted to $V$. By construction, $V$ contains $a_{k}$ and $b_{k}$ for $k$ big enough, which have the same image by $f$. This contradicts the injectivity of $f$ in $V$. ∎
Lemma 2.8 ($p$-adic closed forms versus $p$-adic exact forms).
Let $\ell$ and $k$ be integers with $0\leqslant k\leqslant\ell$ and $\ell\geqslant 1$. Let $p$ be a prime number. Let $M$ be an $\ell$-dimensional $p$-adic analytic manifold and let $Q$ be a compact submanifold of $M$. Let $\alpha$ be a closed $p$-adic analytic $k$-form on $M$ such that
$\alpha_{m}=0\quad\forall m\in Q.$
Then, there is a $p$-adic tubular neighborhood $U$ of $Q$ such that $\alpha$ is an exact $p$-adic analytic form when restricted to $U$.
Proof.
By Lemma 2.7, there exists a $p$-adic analytic tubular neighborhood $V$ of $Q$ as in Definition 2.6. This means that there exists a $p$-adic analytic diffeomorphism $f:N_{k}\to V$, where
$\mathrm{N}(Q)_{k}=\Big\{(x,v)\in\mathrm{T}M:\,\,v^{\mathrm{T}}u=0\text{ for all }u\in\mathrm{T}_{x}Q,\,\,\|v\|_{p}\leqslant p^{-k}\Big\}.$
For any $t\in\mathbb{Z}_{p}$, we define $\psi_{t}:V\to V$ by
$\psi_{t}(f(x,v))=f(x,tv).$
This is well defined because $f$ is a $p$-adic analytic diffeomorphism, and by definition is a $p$-adic analytic family of maps. Note that $\psi_{0}(x)\in Q$ and $\psi_{1}(x)=x$. Now we have, by equations (A.3) and (A.1), that
(2.7) $\frac{\mathrm{d}}{\mathrm{d}t}\psi_{t}^{*}\alpha=\psi_{t}^{*}\mathrm{L}_{\dot{\psi}_{t}}\alpha=\psi_{t}^{*}(\imath_{\dot{\psi}_{t}}\mathrm{d}\alpha+\mathrm{d}\imath_{\dot{\psi}_{t}}\alpha)=\psi_{t}^{*}\mathrm{d}\imath_{\dot{\psi}_{t}}\alpha=\mathrm{d}\psi_{t}^{*}\imath_{\dot{\psi}_{t}}\alpha.$
Let $\beta_{t}=\psi_{t}^{*}\imath_{\dot{\psi}_{t}}\alpha$. Equation (2.7) implies that
$\frac{\mathrm{d}}{\mathrm{d}t}\psi_{t}^{*}\alpha=\mathrm{d}\beta_{t}.$
Since $\beta_{t}$ is a $p$-adic analytic family of (k1)(k-1)-forms, there exists another $p$-adic analytic family of (k1)(k-1)-forms $\gamma_{t}$ such that $\gamma_{0}=0$ and $\frac{\mathrm{d}}{\mathrm{d}t}\gamma_{t}=\beta_{t}$, and $\frac{\mathrm{d}}{\mathrm{d}t}\psi_{t}^{*}\alpha=\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{d}\gamma_{t}$. Since both sides are analytic on $t$, and for $t=0$ we have that $\psi_{0}^{*}\alpha=0=\mathrm{d}\gamma_{0}$ (because $\psi_{0}(x)\in Q$ and $\alpha$ vanishes in $Q$), for some small $t_{0}$ we have that
$\psi_{t_{0}}^{*}\alpha=\mathrm{d}\gamma_{t_{0}}.$
Let $U\subset V$ be the image of $\psi_{t_{0}}$. Then $\alpha|_{U}=\mathrm{d}((\psi_{t_{0}}^{-1})^{*}\gamma_{t_{0}}),$ as we wanted. ∎
The last two results in this section concern symplectomorphisms between unions of balls.
Lemma 2.9.
Let $p$ be a prime number. Let $r$ be a $p$-adic absolute value. The $2n$-dimensional $p$-adic ball of radius $pr$ is a disjoint union of $p^{2n}$ $2n$-dimensional $p$-adic balls of radius $r$.
Proof.
If $r=p^{-k}$ and $m$ is a point in $\mathrm{B}(0,pr)=\mathrm{B}(0,p^{1-k})$, then the digits at the right of the position k1k-1 in all coordinates of $m$ are 0. We can divide the ball in $p^{2n}$ balls according to the digit at position $k$ of each one of the $2n$ coordinates; these balls are a translation of $\mathrm{B}(0,p^{-k})$, as we wanted. ∎
Lemma 2.10.
Let $p$ be a prime number. Let $B,B_{1},\ldots,B_{k}$ be $p$-adic balls such that $\operatorname{vol}(B)=\sum_{i=1}^{k}\operatorname{vol}(B_{i})$, where $\operatorname{vol}$ denotes $p$-adic volume. Then $B$ is symplectomorphic to $\bigcup_{i=1}^{k}B_{i}$.
Proof.
Let $v$ be the smallest volume among the balls $B_{i}$. If $\operatorname{vol}(B)=v$, there is nothing to prove. Otherwise, $\operatorname{vol}(B)$ is a multiple of $p^{2n}v$ and the same happens with all volumes greater than $v$. Hence, there are at least $p^{2n}$ balls with volume $v$. Using Lemma 2.9, we can unify these $p^{2n}$ balls in one with the immediately higher radius. Now we repeat the operation until there is only one ball. ∎
3. Generalization of Moser’s Path Method to $p$-adic geometry: proof of Theorem A
Now we use one of the previous results to prove the $p$-adic analytic Moser’s Path Method (Theorem A). This method requires additional hypotheses when compared to the real case: not only we need the vector field $X_{t}$ to vanish at the compact submanifold, but also the partial derivatives of its components. Furthermore, we are restricting to the case when $X_{t}$ is given by a power series.
Proof of Theorem A
We consider the flow of a point $m^{\prime}$ in $M$ by the vector field $X_{t}$. In terms of local coordinates centered at $m\in Q$, this implies solving an initial value of the form (2.1), where:
  • the variable $x$ is $t$,
  • the variables $y_{1},\ldots,y_{\ell}$ are the local coordinates,
  • $x_{0}=0$,
  • $U_{1}=\mathbb{Z}_{p}$,
  • $U_{2}=p^{-d}\mathbb{Z}_{p}$,
  • $v=0$,
  • the initial values are the coordinates of $m^{\prime}$, and
  • the series $f_{i}$, $i\in\{1,\ldots,\ell\}$, correspond to the components of $X_{t}$.
The series $f_{i}$, $i\in\{1,\ldots,\ell\}$, satisfy the conditions of Lemma 2.1, hence there exists an open neighborhood $U_{m}$ of $m$ such that, if $m^{\prime}\in U_{m}$, it has a unique flow determined by a power series in $t$ which converges in $\mathbb{Z}_{p}$. Let
$U=\bigcup_{m\in Q}U_{m}$
and let $\psi_{t}$ be this flow.
Then (see (A.2)) we have that
$\frac{\mathrm{d}}{\mathrm{d}t}\psi_{t}^{*}\omega_{t}=\psi_{t}^{*}\mathrm{L}_{X_{t}}\omega_{t}+\psi_{t}^{*}\frac{\mathrm{d}}{\mathrm{d}t}\omega_{t}=\psi_{t}^{*}\left(\frac{\mathrm{d}}{\mathrm{d}t}\omega_{t}+\mathrm{L}_{X_{t}}\omega_{t}\right)=0.$
The components of $\psi_{t}^{*}\omega_{t}$ are power series in $t$. If their derivative is 0, then they must be constant. Since for $t=0$ this equals $\omega_{0}$, this must hold for all $t\in\mathbb{Z}_{p}$, and we are done.
4. The Darboux theorem on $p$-adic analytic symplectic manifolds: proof of Theorem B
Now we prove Darboux’s Theorem in the $p$-adic analytic case. This uses the $p$-adic Moser’s Path Method (Theorem A), as well as two of the other results in Section 2.
Theorem 4.1.
Let $n$ be a positive integer. Let $p$ be a prime number. Let $M$ be a $2n$-dimensional $p$-adic analytic manifold. Let $\omega_{0}$ and $\omega_{1}$ be two $p$-adic analytic symplectic forms on $M$. Let $m\in M$. Then there exist open neighborhoods $U_{0},U_{1}$ of $m$ and a diffeomorphism $\psi:U_{0}\to U_{1}$ such that $\psi^{*}(\omega_{1}|_{U_{1}})=\omega_{0}|_{U_{0}}$.
Proof.
Step 1: find region where the difference is exact. By [9, Theorem 9.3] any two $p$-adic linear symplectic forms are always linearly symplectomorphic, hence we may assume that
$(\omega_{0})_{m}=(\omega_{1})_{m}.$
Let
$\alpha=\omega_{1}-\omega_{0}.$
Then $\alpha$ is a closed $2$-form and $\alpha_{m}=0$. By Lemma 2.8, there exists an open subset $U\subset M$ such that $m\in U$ and $\alpha$ is exact when restricted to $U$. This means $\alpha=\mathrm{d}\beta$ for some $p$-adic analytic $1$-form $\beta$ in $U$.
Step 2: find region where the forms are globally analytic. Let $(x_{1},\ldots,x_{2n})$ be local coordinates around $m$. Since $\omega_{0}$ and $\beta$ are analytic, there exists an open subset $U_{2}\subset U$ such that $m\in U_{2}$ and the components of $\omega_{0}$ and $\beta$ are given by power series in the local coordinates $(x_{1},\ldots,x_{2n})$. This implies the same for $\alpha$, because $\alpha=\mathrm{d}\beta$.
Step 3: decompose $\beta$. We now write
$\beta=\beta_{1}+\beta_{2},$
where the power series of $\beta_{1}$ contains only terms of degree 0 and $1$, and that of $\beta_{2}$ contains only terms of higher degree. Hence the power series of $\mathrm{d}\beta_{2}$ (in the same coordinates) contains only non-constant terms, and $(\mathrm{d}\beta_{2})_{m}=0$. Hence
$(\mathrm{d}\beta_{1})_{m}=(\mathrm{d}\beta_{1}+\mathrm{d}\beta_{2})_{m}=(\mathrm{d}\beta)_{m}=\alpha_{m}=0.$
But $\mathrm{d}\beta_{1}$ must be a constant, so $\mathrm{d}\beta_{1}=0$, and $\mathrm{d}\beta_{2}=\mathrm{d}\beta=\alpha$.
Step 4: shrink the region. Let
$\Omega_{0},A\in\mathcal{M}_{2n}(\mathbb{Q}_{p}[[x_{1},\ldots,x_{2n}]])$
be matrices such that
$(\omega_{0})_{x}=\sum_{i=1}^{2n}\sum_{j=1}^{2n}(\Omega_{0})_{ij}(x)\,\mathrm{d}x_{i}\wedge\mathrm{d}x_{j}$
and
$\alpha_{x}=\sum_{i=1}^{2n}\sum_{j=1}^{2n}A_{ij}(x)\,\mathrm{d}x_{i}\wedge\mathrm{d}x_{j}.$
Let $d=2$ if $p=2$ and otherwise $d=1$. Let $T$ be the ball in $\overline{\mathbb{Q}_{p}}$ centered at 0 with radius $p^{d}$. We define
$\chi:U_{2}\times T$ $\to\mathbb{Q}_{p}$
$(x,t)$ $\mapsto\det(\Omega_{0}(x)+tA(x)).$
$\chi$ is a continuous function because the determinant of a matrix is a continuous function on its entries and the entries of $\Omega_{0}(x)$ and $A(x)$ vary continuously with $x$, because $\omega_{0}$ and $\alpha$ are $p$-adic analytic forms. This implies that $S=\chi^{-1}(0)$ is a closed subset of $U_{2}\times T$. Moreover, $S$ does not contain points with $x=m$, because
$\chi(m,t)=\det\Big(\Omega_{0}(m)+tA(m)\Big)=\det(\Omega_{0}(m))\neq 0.$
This implies that there exists an open subset $U_{3}\subset U_{2}$ with $m\in U_{3}$ such that $S\cap(U_{3}\times T)=\varnothing.$ This means that, for all $(x,t)\in U_{3}\times T$,
$\det(\Omega_{0}(x)+tA(x))\neq 0.$
Step 5: define the vector field. Let
$b\in(\mathbb{Q}_{p}[[x_{1},\ldots,x_{2n}]])^{2n}$
be the vector of components of $\beta_{2}$ in local coordinates. For $t\in T$, let $X_{t}$ be the vector field on $U_{3}$ such that
(4.1) $X_{t}(x)=-\Big(\Omega_{0}(x)+tA(x)\Big)^{-1}b(x).$
Then (4.1) gives a $p$-adic analytic family of vector fields, and it satisfies that $\imath_{X_{t}}(\omega_{0}+t\alpha)=-\beta_{2}$ for all $t\in T$. Hence, by (A.1),
$\frac{\mathrm{d}}{\mathrm{d}t}(\omega_{0}+t\alpha)+\mathrm{L}_{X_{t}}(\omega_{0}+t\alpha)=\alpha+\imath_{X_{t}}\mathrm{d}(\omega_{0}+t\alpha)+\mathrm{d}\imath_{X_{t}}(\omega_{0}+t\alpha)=\alpha+0-\mathrm{d}\beta_{2}=0.$
Step 6: apply Moser’s trick. Now we check the hypothesis of Theorem A:
  • The components of $X_{t}$ are rational functions in $t$ which are defined for all $t\in T$. By Lemma 2.4, the power series of $X_{t}$ converges in $\mathrm{B}(0,p^{d})=p^{-d}\mathbb{Z}_{p}$.
  • The power series $b(x)$ is that of $\beta_{2}$, which has no terms of degree 0 or $1$. By equation (4.1), the same holds for $X_{t}(x)$ in the variables of $x$.
Thus we can apply Theorem A and obtain an open subset $U_{0}\subset U_{3}$ and $\psi_{t}:U_{0}\to M$ $p$-adic analytic and injective such that $\psi_{t}^{*}(\omega_{0}+t\alpha)=\omega_{0}|_{U_{0}}$. This in particular holds for $t=1$, and $\psi_{1}^{*}\omega_{1}=\psi_{1}^{*}(\omega_{0}+\alpha)=\omega_{0}|_{U_{0}}$. We conclude the proof by taking $U_{1}=\psi_{1}(U_{0})$ and $\psi=\psi_{1}$. See Figure 3 for an illustration of the proof. ∎
Figure 3. Illustration of the proof of Theorem 4.1. First figure: the form $\alpha=\omega_{1}-\omega_{0}$ is exact on $U$, that is, it is $\mathrm{d}\beta$ for some 11-form $\beta$, and this form is given by a power series in a subset $U_{2}\subset U$. Second figure: we determine $S$ as a subset of $U_{2}\times T$ and a subset $U_{3}\subset U_{2}$ that avoids the points in $S$. Third figure: the vector field $X_{t}$ on $U_{3}$. Fourth figure: the resulting sets $U_{0}$ and $U_{1}$.
Proof of Theorem B
This follows from Theorem 4.1 taking $\omega_{0}=\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}$.
5. The Darboux-Weinstein theorem on $p$-adic analytic symplectic manifolds: proof of Theorem C
Now we prove another classical result about the local behavior of symplectic forms. This is similar to Darboux’s Theorem, but the starting point is given by two symplectic forms coinciding on a compact submanifold. The idea of the proof is as follows: like in the real case, we want to apply Moser’s Path Method, but the $p$-adic equivalent of this result (Theorem A) has additional hypotheses: not only it requires that the vector field vanishes at the compact submanifold, but we also need its derivatives to vanish, and it must be given by a power series. This means that we need to divide the tubular neighborhood given by Lemma 2.8 in several parts, and use local coordinates in each part which allow us to attain the hypotheses of the $p$-adic Moser’s Path Method.
Proof of Theorem C
The proof we carry out next is a generalization of the proof of Theorem 4.1.
Step 1: find tubular neighborhood. Let
$\alpha=\omega_{1}-\omega_{0}.$
Then $\alpha$ is a closed $2$-form and $\alpha_{m}=0$ for all $m\in Q$. By Lemma 2.8, there exists a tubular neighborhood $U$ of $Q$ such that $\alpha$ is exact when restricted to $U$. This means $\alpha=\mathrm{d}\beta$ for some $p$-adic analytic $1$-form $\beta$ in $U$.
Let $q$ be the dimension of $Q$. Since $U$ is a tubular neighborhood of $Q$, each point in $U$ can be identified with a point in $\mathrm{N}(Q)_{k}$ for some $k\in\mathbb{N}$, which is $(x,v)$ where $x\in Q$ and $v\in(p^{k}\mathbb{Z}_{p})^{2n-q}.$
Step 2: find open subsets of $U$. Since $\omega_{0}$ and $\beta$ are analytic, we can decompose $U$ in a family $\{U_{i}\}_{i\in I}$ of pairwise disjoint sets such that $\omega_{0}$ and $\beta$ are given by power series in some local coordinates in each part. This implies the same for $\alpha$, because $\alpha=\mathrm{d}\beta$. Since $Q$ is compact, we can extract a finite subfamily $\{U_{i}\}_{i\in\{1,\ldots,s\}}$ which covers $Q$; we assume that each $U_{i}$ intersects $Q$, because otherwise it can be deleted from the subfamily.
For $i\in\{1,\ldots,s\}$, let
$Q_{i}=Q\cap U_{i}.$
We can choose the local coordinates in $U_{i}$ as $(x_{1},\ldots,x_{2n})$ such that
$Q_{i}=\Big\{x\in U_{i}:x_{q+1}=\ldots=x_{2n}=0\Big\}.$
Step 3: decompose $\beta$. We now write
$\beta=\beta_{i}+\beta_{i}^{\prime},$
where $\beta_{i}$ and $\beta_{i}^{\prime}$ are analytic $1$-forms in $U_{i}$, the power series of $\beta_{i}^{\prime}$ contains only terms of degree 0 and $1$ in the last 2nq2n-q local coordinates $x_{q+1},\ldots,x_{2n}$, and that of $\beta_{i}$ contains only terms of higher degree in those coordinates. This implies that the power series of $\mathrm{d}\beta_{i}$ (in the same coordinates) contains only terms involving $x_{q+1},\ldots,x_{2n}$. Hence, for all $m\in Q$, $(\mathrm{d}\beta_{i})_{m}=0$ and
$(\mathrm{d}\beta_{i}^{\prime})_{m}=(\mathrm{d}\beta_{i}+\mathrm{d}\beta_{i}^{\prime})_{m}=(\mathrm{d}\beta)_{m}=\alpha_{m}=0.$
This means that $\mathrm{d}\beta_{i}^{\prime}$ contains only terms involving $x_{q+1},\ldots,x_{2n}$. But these terms are derivatives of terms in $\beta_{i}^{\prime}$, and every term in $\beta_{i}^{\prime}$ has degree 0 or $1$ in those variables, which means that its derivative, respect to any variable, will not involve them. Hence $\mathrm{d}\beta_{i}^{\prime}$ must be constantly 0, and
$\mathrm{d}\beta_{i}=\mathrm{d}\beta=\alpha.$
Step 4: reduce the sets $U_{i}$. Let
$\Omega_{0i},A_{i}\in\mathcal{M}_{2n}(\mathbb{Q}_{p}[[x_{1},\ldots,x_{2n}]])$
be matrices such that
$(\omega_{0})_{x}=\sum_{k=1}^{2n}\sum_{j=1}^{2n}(\Omega_{0i})_{kj}(x)\,\mathrm{d}x_{k}\wedge\mathrm{d}x_{j}$
and
$\alpha_{x}=\sum_{k=1}^{2n}\sum_{j=1}^{2n}A_{kj}(x)\,\mathrm{d}x_{k}\wedge\mathrm{d}x_{j}.$
Let $d=2$ if $p=2$ and otherwise $d=1$. Let $T$ be the ball in $\overline{\mathbb{Q}_{p}}$ centered at 0 with radius $p^{d}$. We define
$\chi_{i}:U_{i}\times T$ $\to\mathbb{Q}_{p}$
$(x,t)$ $\mapsto\det(\Omega_{0i}(x)+tA_{i}(x)).$
Similarly to the proof of Theorem 4.1, the function $\chi_{i}$ is continuous because the determinant of a matrix is a continuous function on its entries and the entries of $\Omega_{0i}(x)$ and $A_{i}(x)$ vary continuously with $x$, because $\omega_{0}$ and $\alpha$ are $p$-adic analytic forms. Hence $S=\chi_{i}^{-1}(0)$ is closed in $U_{i}\times T$. Furthermore, $S$ does not contain points with $x\in Q$, because if $x\in Q$
$\chi_{i}(x,t)=\det\Big(\Omega_{0i}(x)+tA_{i}(x)\Big)=\det(\Omega_{0i}(x))\neq 0.$
This implies that there exists an open subset $U_{i}^{\prime}\subset U_{i}$ with $Q_{i}\subset U_{i}^{\prime}$ such that $S\cap(U_{i}^{\prime}\times T)=\varnothing.$ This means that, for all $(x,t)\in U_{i}^{\prime}\times T$,
$\det(\Omega_{0i}(x)+tA_{i}(x))\neq 0.$
Step 5: find the vector field. Let
$b_{i}\in(\mathbb{Q}_{p}[[x_{1},\ldots,x_{2n}]])^{2n}$
be the vector of components of $\beta_{i}$ in local coordinates. For $t\in T$, let $X_{i,t}$ be the vector field on $U_{i}^{\prime}$ such that
(5.1) $X_{i,t}(x)=-\Big(\Omega_{0i}(x)+tA_{i}(x)\Big)^{-1}b_{i}(x).$
This is a $p$-adic analytic family of vector fields, and it satisfies that $\imath_{X_{i,t}}(\omega_{0}+t\alpha)=-\beta_{i}$ for all $t\in T$, which implies
$\frac{\mathrm{d}}{\mathrm{d}t}(\omega_{0}+t\alpha)+\mathrm{L}_{X_{i,t}}(\omega_{0}+t\alpha)=\alpha+\imath_{X_{i,t}}\mathrm{d}(\omega_{0}+t\alpha)+\mathrm{d}\imath_{X_{i,t}}(\omega_{0}+t\alpha)=\alpha+0-\mathrm{d}\beta_{i}=0.$
Step 6: apply Moser’s trick. Now we check the hypothesis of Theorem A:
  • The components of $X_{i,t}$ are rational functions in $t$ which are defined for all $t\in T$. By Lemma 2.4, the power series of $X_{i,t}$ converges in $\mathrm{B}(0,p^{d})=p^{-d}\mathbb{Z}_{p}$.
  • The power series $b_{i}(x)$ is that of $\beta_{i}$, which has no terms of degree 0 or $1$ in the variables $x_{q+1},\ldots,x_{2n}$. By equation (5.1), the same holds for $X_{t}(x)$ in the variables of $x$. If we change the origin of the local coordinates to another point in $Q$, we are only changing the variables $x_{1},\ldots,x_{q}$, so the same assertion still holds. This implies that there are no terms of total degree 0 or $1$, regardless of which point in $Q$ we are taking as the origin of the coordinates.
Thus we can apply Theorem A and obtain an open subset $U_{i}^{\prime\prime}\subset U_{i}^{\prime}$ and $\psi_{i,t}:U_{i}^{\prime\prime}\to M$ analytic which fixes $Q$ and such that
$\psi_{i,t}^{*}(\omega_{0}+t\alpha)=\omega_{0}|_{U_{i}^{\prime\prime}}.$
This in particular holds for $t=1$, and
$\psi_{i,1}^{*}\omega_{1}=\psi_{i,1}^{*}(\omega_{0}+\alpha)=\omega_{0}|_{U_{i}^{\prime\prime}}.$
Since the $U_{i}$’s are disjoint, we can now define
$U_{0}=\bigcup_{i=1}^{s}U_{i}^{\prime\prime}$
and
$\psi(x)=\psi_{i,1}(x)$
if $x\in U_{i}^{\prime\prime}$. This satisfies that $\psi^{*}\omega_{1}|_{\psi(U_{0})}=\omega_{0}|_{U_{0}}$, and the proof is complete. See Figure 4 for an illustration of the proof.
Figure 4. Steps of the proof of Theorem C. First figure: the manifold $M$, the compact submanifold $Q$, and a tubular neighborhood $U$ of $Q$. Second figure: open subsets $U_{1},U_{2},U_{3},U_{4}$ covering $Q$ such that $\omega_{0}$ and $\beta$ are given as power series. Third figure: we reduce $U_{i}$ to $U_{i}^{\prime}$ to avoid the points in $S_{i}$. Fourth figure: the vector field $X_{i,t}$ in $U_{i}^{\prime}$ and the resulting set $U_{i}^{\prime\prime}$. Fifth figure: the set $U_{0}$ is the union of all $U_{i}^{\prime\prime}$.
6. Global classification of $p$-adic analytic symplectic manifolds: proof of Theorems D and E
In this section we prove our classification of $p$-adic analytic symplectic manifolds. We start with Theorem D about the possible classes.
Proof of Theorem D
  1. (1)
    It is enough to show that $\operatorname{vol}(\mathcal{M}_{d,v})=v$. This happens because
    $\operatorname{vol}(\mathcal{M}_{d,v})=\operatorname{vol}\left(\bigcup_{k=-\infty}^{m}\bigcup_{i=1}^{a_{k}}\mathrm{B}_{ki}\right)=\sum_{k=-\infty}^{m}\sum_{i=1}^{a_{k}}p^{dk}=\sum_{k=-\infty}^{m}a_{k}p^{dk}=v.$
  2. (2)
    It is enough to show that $\operatorname{vol}(\mathcal{M}_{d,v}^{\prime})=v$. This happens because
    $\operatorname{vol}(\mathcal{M}_{d,v}^{\prime})=\operatorname{vol}\left(\bigcup_{k=\ell}^{m}\bigcup_{i=1}^{a_{k}}\mathrm{B}_{ki}\right)=\sum_{k=\ell}^{m}\sum_{i=1}^{a_{k}}p^{dk}=\sum_{k=\ell}^{m}a_{k}p^{dk}=v.$
Figure 5. The balls $\mathrm{B}_{ki}$, for $p=3$, $d=2$, $0\leqslant k\leqslant 2$, and $1\leqslant i\leqslant 8$. The ball $\mathrm{B}_{ki}$ has radius $p^{k}$ and a color which depends on $i$.
In order to prove Theorem E, the starting point is to use Theorem B to write a $p$-adic analytic symplectic manifold as a disjoint union of balls.
Lemma 6.1.
Let $p$ be a prime number. Let $(M,\omega)$ be a second countable $p$-adic analytic symplectic manifold. Then $(M,\omega)$ is symplectomorphic to a disjoint union of $p$-adic balls, each one endowed with the standard symplectic form $\omega_{\mathrm{std}}$.
Proof.
By Theorem B, for each point $m\in M$, there exists an open subset $U_{m}$ of $M$ such that $m\in M$ and $U_{m}$ is symplectomorphic to an open subset $V_{m}\subset(\mathbb{Q}_{p})^{2n}$ with the standard symplectic form. Since $M$ is a second countable $p$-adic manifold, it is strictly paracompact by [38, Proposition 8.7], which means that the covering $\{U_{m}\}_{m\in M}$ can be refined to one consisting of disjoint sets. Let $\{U^{\prime}_{i}\}_{i\in I}$ be this covering. For each $i\in I$, there exists $m$ such that $U^{\prime}_{i}\subset U_{m}$. Let $V^{\prime}_{i}$ be the image of $U^{\prime}_{i}$ by the symplectomorphism $U_{m}\to V_{m}$. We have that $V^{\prime}_{i}$ is an open subset of $(\mathbb{Q}_{p})^{2n}$. Hence, it is a disjoint union of $p$-adic balls, and $U^{\prime}_{i}$ is symplectomorphic to that union. Since $M$ is the disjoint union of the $U^{\prime}_{i}$, it is also symplectomorphic to a disjoint union of $p$-adic balls. ∎
This result, together with the last two results in Section 2, will allow us to prove Theorem E.
Proof of Theorem E
In the three cases, by Lemma 6.1, we can suppose that $M$ is a disjoint union of $p$-adic balls. Let $v$ be the volume of $M$.
  1. (1a)
    Since $M$ is noncompact, the union must be infinite, and it is countable because $M$ is second-countable. By Lemma 2.9, we may assume, without loss of generality, that the radii of the balls is less or equal than $1$ (otherwise, we divide all balls of radius greater than $1$ into smaller balls). We now modify the union so that it contains a finite number of balls of each possible radius. In order to achieve this, if there are infinitely many balls of radius $r$, we leave the first as such, decompose the second into $p^{2n}$ balls of radius $r/p$, the third into $p^{4n}$ balls of radius $r/p^{2}$, and so on.
    Let $B_{1},B_{2},\ldots$ be the balls in $M$ with radii $r_{1},r_{2},\ldots$ in decreasing order, that is, with $r_{i}\geqslant r_{i+1}$. We define $a_{0}=0$ and, for each $k\in\mathbb{N}$,
    $a_{k}=\sum_{i=1}^{k}r_{i}^{2n}$
    Since $r_{i}$ and $r_{i+1}$ are $p$-adic absolute values and $r_{i}\geqslant r_{i+1}$, we have $r_{i}/r_{i+1}\in\mathbb{Z}$. This in turn implies that $r_{i}/r_{j}\in\mathbb{Z}$ for all $i,j\in\mathbb{N}$ with $i<j$ and
    $\frac{a_{k}}{r_{k}^{2n}}=\sum_{i=1}^{k}\left(\frac{r_{i}}{r_{k}}\right)^{2n}\in\mathbb{Z}.$
    We also have that
    $\lim_{k\to\infty}a_{k}=\sum_{i=1}^{\infty}r_{i}^{2n}=\operatorname{vol}(M)=\infty.$
    We claim that, for all $m\in\mathbb{N}$, there is $k(m)\in\mathbb{N}$ such that $a_{k(m)}=m$. Suppose otherwise that $a_{k}<m<a_{k+1}$ for some $k\in\mathbb{N}$. Then we have that
    $\frac{a_{k}}{r_{k+1}^{2n}}<\frac{m}{r_{k+1}^{2n}}<\frac{a_{k+1}}{r_{k+1}^{2n}}=\frac{a_{k}}{r_{k+1}^{2n}}+1,$
    which is impossible because all the terms are integers.
    Now the sum of volumes of the balls $B_{k(m)+1},B_{k(m)+2},\ldots,B_{k(m+1)}$ is $a_{k(m+1)}-a_{k(m)}=1$, and by Lemma 2.10, their union is symplectomorphic to a ball of radius $1$. Since $\mathbb{Q}_{p}$ is a countable union of balls of radius $1$, the union of all balls is symplectomorphic to $\mathbb{Q}_{p}$ and we are done.
  2. (1b)
    Again, since $M$ is noncompact, the union must be infinite, and it is countable because $M$ is second-countable. Let $\mathrm{B}_{mi}$ be a ball contained in $\mathcal{M}_{2n,v}$ such that $k$ is the maximum possible. We claim that there exists a subfamily of balls contained in $M$ which is symplectomorphic to $\mathrm{B}_{mi}$.
    Let $B_{1},B_{2},\ldots$ be the balls in $M$ with radii $r_{1},r_{2},\ldots$ in decreasing order, that is, with $r_{i}\geqslant r_{i+1}$. We define $a_{0}=0$ and, for each $k\in\mathbb{N}$,
    $a_{k}=\sum_{i=1}^{k}r_{i}^{2n}$
    Since $r_{i}$ and $r_{i+1}$ are $p$-adic absolute values and $r_{i}\geqslant r_{i+1}$, we have $r_{i}/r_{i+1}\in\mathbb{Z}$. This in turn implies that $r_{i}/r_{j}\in\mathbb{Z}$ for all $i,j\in\mathbb{N}$ with $i<j$ and
    $\frac{a_{k}}{r_{k}^{2n}}=\sum_{i=1}^{k}\left(\frac{r_{i}}{r_{k}}\right)^{2n}\in\mathbb{Z}.$
    We also have that
    $\lim_{k\to\infty}a_{k}=\sum_{i=1}^{\infty}r_{i}^{2n}=\operatorname{vol}(M)=v.$
    Let $k$ be such that $a_{k-1}<p^{2nm}\leqslant a_{k}$. This $k$ exists since $v=\operatorname{vol}(\mathcal{M}_{2n,v})>\operatorname{vol}(\mathrm{B}_{mi})=p^{2nm}$. If $a_{k}=p^{2nm}$, Lemma 2.10 means that the union $B_{1}\cup\ldots\cup B_{k}$ is symplectomorphic to $\mathrm{B}_{mi}$. Suppose otherwise that $a_{k}>p^{2nm}$. Because of our choice of $m$, $p^{2n(m+1)}>v$, hence $r_{k}^{2n}<p^{2n(m+1)}$ and $r_{k}<p^{m+1}$, which implies $r_{k}<p^{m}$. Then we have that
    $\frac{a_{k-1}}{r_{k}^{2n}}<\frac{p^{2nm}}{r_{k}^{2n}}<\frac{a_{k}}{r_{k}^{2n}}=\frac{a_{k-1}}{r_{k}^{2n}}+1,$
    which is impossible because all the terms are integers.
    We have obtained so far a symplectic embedding from $\mathrm{B}_{mi}$ to $M$ whose image is a union of a subfamily of balls. By removing from $M$ these balls and iterating the construction, we can obtain a symplectic embedding $\varphi$ from $\mathcal{M}_{2n,v}$ to $M$ whose image is the union of a subfamily of balls. Suppose that $\varphi$ is not a symplectomorphism. Then there is a ball $B_{0}$ in $M$ which is not in the image of $\varphi$. This means that
    $v=\operatorname{vol}(M)\geqslant\operatorname{vol}(\mathcal{M}_{2n,v})+\operatorname{vol}(B_{0})=v+\operatorname{vol}(B_{0})>v,$
    which is a contradiction. Hence, $\varphi$ is a symplectomorphism and we are done.
  3. (1c)
    In this case the union is finite. Let $r$ be the smallest radius of all balls in $M$. Since $v$ is a multiple of $r$, by the definition of $\mathcal{M}_{2n,v}^{\prime}$, all balls in $\mathcal{M}_{2n,v}^{\prime}$ have at least radius $r$. By Lemma 2.9, $M$ and $\mathcal{M}_{2n,v}^{\prime}$ can be written as disjoint unions of balls of radius $r$. Since the two manifolds have the same volume, the number of balls in both unions must be the same, and $M$ and $\mathcal{M}_{2n,v}^{\prime}$ are symplectomorphic.
  4. (2)
    This is a consequence of the fact that all symplectomorphisms preserve $p$-adic volume.
See Figure 6 for an example of the proof.
Figure 6. An example of the proof of Theorem E. The colored balls form a $p$-adic analytic symplectic manifold in $(\mathbb{Q}_{3})^{2}$. Those with radius 33 can be mapped to $\mathrm{B}_{11}$, $\mathrm{B}_{12}$, $\mathrm{B}_{13}$ and $\mathrm{B}_{14}$ according to their colors, while those of radius 11 are mapped to $\mathrm{B}_{01}$, $\mathrm{B}_{02}$, $\mathrm{B}_{03}$, $\mathrm{B}_{04}$, $\mathrm{B}_{05}$, and the remaining nine to $\mathrm{B}_{15}$.
7. Application to nonlinear Schrödinger equation
We will now give an explicit example of the $p$-adic Darboux’s Theorem based on the ideas for the real case in [5].
The Discrete Nonlinear Schrödinger Model (dNLS) is a Hamiltonian dynamical system given by
$H(\psi)=\sum_{j\in J}\left(\epsilon(\psi_{j}\bar{\psi}_{j+1}+\bar{\psi}_{j}\psi_{j+1})+\frac{\gamma}{2}|\psi_{j}|^{4}\right)$
where $J=\{1,\ldots,n\}$ or $J=\mathbb{Z}$, $\epsilon\in\mathbb{R}^{+}$, $\gamma\in\mathbb{R}$, and $\psi_{j}\in\mathbb{C}$ for each $j\in J$. The standard symplectic form is given in those coordinates by
$\omega_{\mathrm{std}}=\mathrm{i}\sum_{j\in J}^{n}\mathrm{d}\psi_{j}\wedge\mathrm{d}\bar{\psi}_{j}$
and the movement equation is
$\mathrm{i}\frac{\mathrm{d}}{\mathrm{d}t}\psi_{j}=\epsilon(\psi_{j+1}+\psi_{j-1})+\gamma|\psi_{j}|^{2}\psi_{j}.$
A generalization of this model changes the symplectic form to
$\omega=\mathrm{i}\sum_{j\in J}\frac{\mathrm{d}\psi_{j}\wedge\mathrm{d}\bar{\psi}_{j}}{1+\mu|\psi_{j}|^{2}}$
and takes as Hamiltonian
$H(\psi)=\sum_{j\in J}\left(\epsilon(\psi_{j}\bar{\psi}_{j+1}+\bar{\psi}_{j}\psi_{j+1})+\frac{\gamma}{\mu}|\psi_{j}|^{2}-\frac{\gamma}{\mu^{2}}\log(1+\mu|\psi_{j}|^{2})\right)$
where $J=\{1,\ldots,n\}$ or $J=\mathbb{Z}$, $\epsilon,\mu\in\mathbb{R}^{+}$, $\gamma\in\mathbb{R}$, and $\psi_{j}\in\mathbb{C}$ for each $j\in J$. Now the movement equation is
$\mathrm{i}\frac{\mathrm{d}}{\mathrm{d}t}\psi_{j}=\epsilon(1+\mu|\psi_{j}|^{2})(\psi_{j+1}+\psi_{j-1})+\gamma|\psi_{j}|^{2}\psi_{j}.$
When $\gamma=0$, it is known as Ablowitz-Ladik Model; when $\gamma\neq 0$, it is the Salerno Model. When $\mu$ tends to 0, we recover the dNLS model. According to [5], after changing coordinates to those given by Darboux’s Theorem, the system also becomes the dNLS model.
For our purposes, we will use real coordinates $(x_{1},y_{1},\ldots,x_{n},y_{n})$ instead of complex $(\psi_{1},\ldots,\psi_{n})$:
$\left\{\begin{aligned} \psi_{j}&=\frac{x_{j}+\mathrm{i}y_{j}}{\sqrt{2}};\\ \bar{\psi}_{j}&=\frac{x_{j}-\mathrm{i}y_{j}}{\sqrt{2}}.\end{aligned}\right$.
We will apply Darboux’s Theorem (Theorem B) to the $p$-adic equivalent of this system, where the $2n$ variables are in $\mathbb{Q}_{p}$ instead of in $\mathbb{R}$. Often it is not possible to obtain an explicit expression of the Darboux coordinates because it involves integrating very complicated expressions, but one can follow the proof of the $p$-adic analytic Darboux’s Theorem and at least derive non-explicitly the $\phi$ and some of its properties. This is the case for the example we are treating in this section.
Though we have a symplectic form on $(\mathbb{Q}_{p})^{2n}$, it is actually the sum of $n$ symplectic forms with the same expression, each one on $(\mathbb{Q}_{p})^{2}$, so we can apply Theorem B to the first two and the conclusion will apply to all of them:
$\omega_{1}(x_{1},y_{1})=\frac{\omega_{\mathrm{std}}(x_{1},y_{1})}{1+\nu(x_{1}^{2}+y_{1}^{2})},$
where $\nu=\mu/2$ and $\omega_{\mathrm{std}}(x_{1},y_{1})=\mathrm{d}x_{1}\wedge\mathrm{d}y_{1}$. We now want to find local coordinates $(x_{0},y_{0})$ around 0 in $(\mathbb{Q}_{p})^{2}$ which are symplectic, that is, in which
$\omega_{1}(x_{1},y_{1})=\omega_{\mathrm{std}}(x_{0},y_{0}).$
Proposition 7.1.
The diffeomorphism corresponding to integrating
$X_{t}=\frac{1+\nu(x_{1}^{2}+y_{1}^{2})}{1+(1-t)\nu(x_{1}^{2}+y_{1}^{2})}\left(1-\frac{\log(1+\nu(x_{1}^{2}+y_{1}^{2}))}{2\nu(x_{1}^{2}+y_{1}^{2})}\right)\begin{pmatrix}x_{1}\\ y_{1}\end{pmatrix}$
takes $\omega_{1}$ to $\mathrm{d}x_{1}\wedge\mathrm{d}y_{1}$ (this is the diffeomorphism $\psi$ found theoretically in Theorem B).
Proof.
First of all, we need to apply a linear transformation to achieve that $\omega_{1}=\omega_{\mathrm{std}}$ at the origin. In this case, this transformation is the identity.
The next step is to make
$\alpha=\omega_{1}-\omega_{\mathrm{std}}=-\frac{\nu(x_{1}^{2}+y_{1}^{2})}{1+\nu(x_{1}^{2}+y_{1}^{2})}\omega_{\mathrm{std}}$
and find a $1$-form $\gamma$ such that $\mathrm{d}\gamma=\alpha$ near the origin. We can write
$\gamma=f(\nu(x_{1}^{2}+y_{1}^{2}))(-y_{1}\mathrm{d}x_{1}+x_{1}\mathrm{d}y_{1})$
so that
$\mathrm{d}\gamma$ $=f(\nu(x_{1}^{2}+y_{1}^{2}))(-\mathrm{d}y_{1}\wedge\mathrm{d}x_{1}+\mathrm{d}x_{1}\wedge\mathrm{d}y_{1})$
$\;\;\;+f^{\prime}(\nu(x_{1}^{2}+y_{1}^{2}))(-2\nu y_{1}^{2}\mathrm{d}y_{1}\wedge\mathrm{d}x_{1}+2\nu x_{1}^{2}\mathrm{d}x_{1}\wedge\mathrm{d}y_{1})$
$=2f(\nu(x_{1}^{2}+y_{1}^{2}))\mathrm{d}x_{1}\wedge\mathrm{d}y_{1}$
$\;\;\;+2f^{\prime}(\nu(x_{1}^{2}+y_{1}^{2}))\nu(x_{1}^{2}+y_{1}^{2})\mathrm{d}x_{1}\wedge\mathrm{d}y_{1},$
and we need
$2f(t)+2tf^{\prime}(t)=-\frac{t}{1+t},$
that is
$2\frac{\mathrm{d}}{\mathrm{d}t}(tf(t))=-\frac{t}{1+t},$
$tf(t)=\int-\frac{t}{2(1+t)}\mathrm{d}t=\frac{1}{2}(\log(1+t)-t),$
and we finally obtain
$f(t)=\frac{1}{2}\left(\frac{\log(1+t)}{t}-1\right).$
Substituting in the expression for $\gamma$,
$\gamma=\left(\frac{\log(1+\nu(x_{1}^{2}+y_{1}^{2}))}{2\nu(x_{1}^{2}+y_{1}^{2})}-1\right)(-y_{1}\mathrm{d}x_{1}+x_{1}\mathrm{d}y_{1}).$
The next step is to obtain the vector field $X_{t}$ such that
$\imath_{X_{t}}(\omega_{\mathrm{std}}+t\alpha)=-\gamma.$
This gives
$\begin{pmatrix}0&-1+t\frac{\nu(x_{1}^{2}+y_{1}^{2})}{1+\nu(x_{1}^{2}+y_{1}^{2})}\\ 1-t\frac{\nu(x_{1}^{2}+y_{1}^{2})}{1+\nu(x_{1}^{2}+y_{1}^{2})}&0\end{pmatrix}X_{t}=\left(1-\frac{\log(1+\nu(x_{1}^{2}+y_{1}^{2}))}{2\nu(x_{1}^{2}+y_{1}^{2})}\right)\begin{pmatrix}-y_{1}\\ x_{1}\end{pmatrix}$
and
$X_{t}$ $=\begin{pmatrix}0&\frac{1+\nu(x_{1}^{2}+y_{1}^{2})}{1+(1-t)\nu(x_{1}^{2}+y_{1}^{2})}\\ -\frac{1+\nu(x_{1}^{2}+y_{1}^{2})}{1+(1-t)\nu(x_{1}^{2}+y_{1}^{2})}&0\end{pmatrix}\left(1-\frac{\log(1+\nu(x_{1}^{2}+y_{1}^{2}))}{2\nu(x_{1}^{2}+y_{1}^{2})}\right)\begin{pmatrix}-y_{1}\\ x_{1}\end{pmatrix}$
$=\frac{1+\nu(x_{1}^{2}+y_{1}^{2})}{1+(1-t)\nu(x_{1}^{2}+y_{1}^{2})}\left(1-\frac{\log(1+\nu(x_{1}^{2}+y_{1}^{2}))}{2\nu(x_{1}^{2}+y_{1}^{2})}\right)\begin{pmatrix}x_{1}\\ y_{1}\end{pmatrix}$.
Now we have $X_{t}$, its integral will give us the required coordinate change. ∎
Remark 7.2.
This vector field can be related to $V_{t}$ in [5, Theorem 2.1] by $X_{t}=-V_{1-t}$. This corresponds exactly to the effect of inverting the coordinate change: we are taking $\omega_{0}$ as the standard form and $\omega_{1}$ as a non-standard form, while they take $\omega_{1}$ as the standard form. This means that the algebraic expression of the vector field in the real and $p$-adic cases is essentially the same. The differences between the proofs of real and $p$-adic Darboux’s theorems are only due to the different topology of both fields.
8. Final remarks and context of the paper
Remark 8.1 (References in $p$-adic/symplectic geometry and context of the paper).
Symplectic geometry originates in the study of planetary motions in the seventeenth and eighteenth centuries and has deep connections with mathematical physics, analysis/PDEs, representation theory and many other subjects (see [16, 31, 32, 37, 46]). For an introduction to symplectic geometry, we recommend the textbooks and surveys [6, 23, 27, 28, 30, 31], and for its connection to the theory of quantum states see [13, 41, 44]. The role of $p$-adic numbers in physics has been extensively studied, see for example [2, 14, 15, 36, 42]. A construction of the $p$-adic symplectic and Heisenberg groups and the Maslov index can be found in [24, 48]. The present paper is part of a larger program proposed by Voevodsky, Warren and the second author in [33] to develop a $p$-adic version of symplectic geometry with the goal of later implementing it using a proof assistant (in the language of homotopy type theory and Voevodsky’s Univalent Foundations [1, 34]). Within this program the results of our paper are probably the most foundational of all: they clarify the local theory completely. In our previous works in $p$-adic symplectic geometry we focused on matrices/integrable systems [8, 9] and flexibility/rigidity [10].
Remark 8.2 (Gromov’s Nonsqueezing).
Theorem E implies that Gromov’s Nonsqueezing Theorem [22] does not hold in general for $p$-adic analytic symplectic manifolds, as we proved explicitly in [10, Theorem B] (see Table 1). Indeed, a $p$-adic cylinder is a noncompact second-countable $2n$-dimensional $p$-adic analytic symplectic manifold with infinite volume, hence Theorem E implies that it is symplectomorphic to $(\mathbb{Q}_{p})^{2n}$.
Remark 8.3 (Second-countable axiom).
Being second-countable is sometimes taken as part of the definition of manifold, see for example [25]. This is often the case of real symplectic geometry: although [28] does not state it explicitly, it can be deduced from its definitions and results that second-countability is being assumed there. In Schneider’s book [38], which concerns $p$-adic analytic manifolds, this is not assumed by default to avoid being too restrictive.
Without the assumption of being second-countable Theorem E is false, because, for example, the union of an uncountable number of copies of $(\mathbb{Q}_{p})^{2n}$ cannot be embedded into $(\mathbb{Q}_{p})^{2n}$, and the statement of Theorem F makes no sense in general, because the volume of a manifold is not necessarily well defined without this hypothesis. Actually, we can construct families, with greater cardinality than the continuum, of (non-second-countable) $p$-adic analytic manifolds which are not even homeomorphic [21]. In view of this, we are tempted to say that the “correct” formulation of $p$-adic symplectic geometry is without assuming second-countability, unlike it is often done in the real case.
Remark 8.4 (Darboux-Weinstein’s Theorem does not follow from Serre’s Theorem).
Theorem C is not a consequence of Theorem F, because the latter ensures the existence of a symplectomorphism but does not impose any condition on it, while the former specifies that the symplectomorphism fixes the compact submanifold $Q$.
Remark 8.5.
In Theorem A, the family of $k$-forms $\omega_{t}$ and the vector field $X_{t}$, seen as power series in the variable $t$, must converge in $p^{-d}\mathbb{Z}_{p}$ in order to apply Moser’s Path Method, while the solution $\psi_{t}$ is only obtained for $t\in\mathbb{Z}_{p}$.
symplectic geometry $p$-adic symplectic geometry
Darboux’s Theorem
Global Darboux’s Theorem (Serre’s classification) ×
Darboux-Weinstein’s Theorem
Gromov’s linear nonsqueezing
Gromov’s nonsqueezing ×
Equivariant Gromov’s nonsqueezing
Few normal forms for integrable systems in dimension 44
(7 forms)
Interesting physical models
Table 1. Comparison of real and $p$-adic symplectic geometry. All the classical results of real symplectic geometry included here also hold in $p$-adic symplectic geometry, with the exceptions of Gromov’s nonsqueezing and the normal forms for integrable systems. In the first case, this is due to the fact that symplectic transformations, in a global and nonlinear sense, have more freedom in the $p$-adic case than in the real case, though locally and linearly they are similar. For the classification of integrable systems, the reason lies behind the higher algebraic complexity of the $p$-adic numbers. See [8, 9, 10] for the proofs.
Appendix A $p$-adic numbers and $p$-adic geometry
A.1. $p$-adic numbers, power series and analytic functions
In this section we explain the concepts we need about the $p$-adic numbers. We recommend Lurie’s lecture [26] and Scholze-Weinstein’s lectures [39] for a general introduction to $p$-adic geometry. The following presentation follows mainly Schneider’s book on $p$-adic Lie groups [38].
Let $p$ be a prime number. For $n\in\mathbb{Z}$, we define the $p$-adic absolute value $|n|_{p}$ of $n$ as
$|n|_{p}=p^{-k},$
where $p^{k}$ is the highest power of $p$ which divides $n$. For $m/n\in\mathbb{Q}$, we define the $p$-adic absolute value of $m/n$ as
$\left|\frac{m}{n}\right|_{p}=\frac{|m|_{p}}{|n|_{p}}.$
We say that $r\in\mathbb{Q}$, with $r\neq 0$, is a $p$-adic absolute value if it is in the image of the $p$-adic absolute value function (equivalently, it is a power of $p$). The field $\mathbb{Q}_{p}$ is the metric completion of $\mathbb{Q}$ with respect to the $p$-adic absolute value.
Let $\ell$ be a positive integer. We define the $p$-adic norm of $(x_{1},\ldots,x_{\ell})\in(\mathbb{Q}_{p})^{\ell}$ by
$\|(x_{1},\ldots,x_{n})\|_{p}=\max\Big\{|x_{i}|_{p}:i\in\{1,\ldots,\ell\}\Big\}.$
Let $x\in(\mathbb{Q}_{p})^{\ell}$ and let $r$ be a $p$-adic absolute value. The ball of center $x$ and radius $r$ in $(\mathbb{Q}_{p})^{\ell}$, denoted by $\mathrm{B}(x,r)$, is the set
$\Big\{y\in(\mathbb{Q}_{p})^{\ell}:\|y-x\|_{p}\leqslant r\Big\}=\Big\{y\in(\mathbb{Q}_{p})^{\ell}:\|y-x\|_{p}<pr\Big\}.$
See Figures 7 and 8 for representations of balls for $p=5$ in dimension $2$ and $3$. A $p$-adic ball is an open and closed set.
Figure 7. A representation of the ball $\mathrm{B}(0,25)$ in $(\mathbb{Q}_{5})^{2}$. Each point is a ball of radius $1$. The red point in the center is $\mathrm{B}(0,1)$ and the green points form $\mathrm{B}(0,5)$.
Figure 8. Left: a representation of the ball $\mathrm{B}(0,25)$ in $(\mathbb{Q}_{5})^{3}$. Each square is a ball of radius $5$. Right: the ball $\mathrm{B}(0,5)$, where each point is a ball of radius 11 and the red point in the center is $\mathrm{B}(0,1)$.
A $p$-adic power series in $(\mathbb{Q}_{p})^{\ell}$ is a series of the form
$\sum_{I=(i_{1},\ldots,i_{\ell})\in\mathbb{N}^{\ell}}a_{I}(x_{1}-x_{01})^{i_{1}}\ldots(x_{\ell}-x_{0\ell})^{i_{\ell}},$
where $(x_{01},\ldots,x_{0\ell})\in(\mathbb{Q}_{p})^{\ell}$ and $a_{I}\in\mathbb{Q}_{p}$. Given a $p$-adic power series $f\in\mathbb{Q}_{p}[[x_{1},\ldots,x_{\ell}]]$, for each $k\in\mathbb{N}$, let $C_{k}$ be the maximum of $|a_{I}|_{p}$ where $I=(i_{1},\ldots,i_{\ell})$ and $\sum_{j=1}^{\ell}i_{j}=k$. Let $r$ be a $p$-adic absolute value. Then it is easy to check that $f$ converges at all points of $\mathrm{B}(x_{0},r)$ if and only if $\lim_{k\to\infty}C_{k}r^{k}=0$. This follows from the fact that a $p$-adic series converges if and only if its terms tend to 0.
Finally, if $U\subset(\mathbb{Q}_{p})^{\ell_{1}}$ and $V\subset(\mathbb{Q}_{p})^{\ell_{2}}$ are open sets, a function $f:U\to V$ is $p$-adic analytic if $U$ can be expressed as $U=\bigcup_{i\in I}U_{i}$ where $U_{i}=\mathrm{B}(x_{i},r_{i})$, for some $x_{i}\in(\mathbb{Q}_{p})^{\ell_{1}}$ and $p$-adic absolute values $r_{i}$, and there are power series $f_{i}$ converging in $U_{i}$ such that $f(x)=f_{i}(x)$ for every $x\in U_{i}$.
The exponential series is defined as
$\exp(x)=\sum_{i=0}^{\infty}\frac{x^{i}}{i!}.$
By [8, Proposition A.11], this series converges in $p\mathbb{Z}_{p}$ if $p\neq 2$, and $4\mathbb{Z}_{2}$ if $p=2$.
A.2. $p$-adic analytic manifolds, vector fields and forms
Let $\ell$ be a positive integer. Let $p$ be a prime number. Let $M$ be a Hausdorff topological space. An $\ell$-dimensional $p$-adic analytic atlas is a set of functions $A=\{\phi:U_{\phi}\to V_{\phi}\}$, where $U_{\phi}\subset M$ and $V_{\phi}\subset(\mathbb{Q}_{p})^{\ell}$ are open subsets, such that:
  • $\phi:U_{\phi}\to V_{\phi}$ is a homeomorphism;
  • for any $\phi,\psi\in A$, the change of charts
    $\psi\circ\phi^{-1}:\phi(U_{\phi}\cap U_{\psi})\to\psi(U_{\phi}\cap U_{\psi})$
    is bi-analytic, i.e. it is analytic with analytic inverse.
A topological space $M$ endowed with an $\ell$-dimensional $p$-adic analytic atlas is called a $p$-adic analytic manifold. The integer $\ell$ is called the dimension of $M$.
Let $M$ and $N$ be $p$-adic analytic manifolds of dimensions $\ell_{1}$ and $\ell_{2}$ respectively. A map $F:M\to N$ is $p$-adic analytic if, for any $m\in M$, there exist neighborhoods $U_{\phi}$ of $m$ and $U_{\psi}$ of $F(m)$ such that $\psi\circ F\circ\phi^{-1}$ is $p$-adic analytic (this composition is a function from a subset of $(\mathbb{Q}_{p})^{\ell_{1}}$ to a subset of $(\mathbb{Q}_{p})^{\ell_{2}}$).
The following concepts are analogous to the real case:
  • Given a $p$-adic analytic manifold $M$, a function $f:M\to\mathbb{Q}_{p}$ is $p$-adic analytic if it is analytic as a map between manifolds. Equivalently, for the charts $\phi$ of $M$, $f|_{U_{\phi}}\circ\phi^{-1}$ is analytic on $\phi(U_{\phi})$. The space of analytic maps $M\to\mathbb{Q}_{p}$ is denoted by $\Omega^{0}(M)$.
  • A tangent vector to $m\in M$ is a linear map $v:\Omega^{0}(M)\to\mathbb{Q}_{p}$ such that $v(fg)=v(f)g(m)+f(m)v(g)$ for all $f,g\in\Omega^{0}(M)$. We denote the space of tangent vectors to $m$ by $\mathrm{T}_{m}M$ and by $\mathrm{T}M=\bigcup_{m\in M}\mathrm{T}_{m}M$ the $p$-adic tangent bundle, which is again defined by analogy with the real case.
[…]
  • A $p$-adic analytic (embedded) submanifold $Q$ of $M$ is a subset of $M$ for each point $m\in Q$ there exists a chart $\phi:U\to V$ such that $m\in U$ and $\phi(U\cap Q)$ is a $k$-slice of $\phi(U)$, that is, an open set of the form
    $\{(x_{1},\ldots,x_{k},x_{k+1},\ldots,x_{n})\in\phi(U):x_{k+1}=c_{k+1},\ldots,x_{n}=c_{n}\},$
    for some constants $c_{k+1},\ldots,c_{n}\in\mathbb{Q}_{p}$.
  • The pullback $F^{*}(\alpha)\in\Omega^{k}(M)$ of a form $\alpha\in\Omega^{k}(N)$ by $F$, the push-forward $F_{*}(v)\in\mathrm{T}_{m}N$ of a vector $v\in\mathrm{T}_{m}M$, and, if $F$ is bi-analytic, the push-forward $F_{*}(X)\in\mathfrak{X}(N)$ of a vector field $X\in\mathfrak{X}(M)$, are defined analogously to the real case.
  • Similarly for the wedge operator, the differential operator $\mathrm{d}$, the contraction operator $\imath_{X}$ and the Lie derivative $\mathrm{L}_{X}$ for a vector field $X$. Like in the real case, Cartan’s Magic Formula holds in the $p$-adic case:
    (A.1) $\mathrm{L}_{X}\omega=\imath_{X}\mathrm{d}\omega+\mathrm{d}\imath_{X}\omega.$
  • Also, if $\psi_{t}$ is the flow of $X_{t}$ and $\{\omega_{t}\}_{t\in p^{-d}\mathbb{Z}_{p}}$ is a $p$-adic analytic family of $p$-adic analytic $k$-forms on $M$, then
    (A.2) $\frac{\mathrm{d}}{\mathrm{d}t}\psi_{t}^{*}\omega_{t}=\psi_{t}^{*}\left(\mathrm{L}_{X_{t}}\omega_{t}+\frac{\mathrm{d}\omega_{t}}{\mathrm{d}t}\right).$
    In particular,
    (A.3) $\frac{\mathrm{d}}{\mathrm{d}t}\psi_{t}^{*}\omega=\psi_{t}^{*}\mathrm{L}_{X_{t}}\omega.$
Let $n$ be a positive integer. Let $p$ be a prime number. Following Pelayo-Voevodsky-Warren [33, Section 7.2], a $2n$-dimensional $p$-adic (analytic) symplectic manifold is a pair $(M,\omega)$, where $M$ is a $2n$-dimensional $p$-adic analytic manifold and $\omega$ is a closed non-degenerate $p$-adic analytic $2$-form, that is, $\mathrm{d}\omega=0$ and for all $m\in M$ and $u\in\mathrm{T}_{m}M,u\neq 0,$ there is $v\in\mathrm{T}_{m}M$ such that $\omega(u,v)\neq 0$. The form $\omega$ is called a $p$-adic analytic symplectic form or simply a $p$-adic symplectic form.
Let $p$ be a prime number. Let $A$, $B$ and $T$ be $p$-adic analytic manifolds. Let $f_{t}:A\to B$ be a $p$-adic analytic map for each $t\in T$. We say that $\{f_{t}\}_{t\in T}$ is a $p$-adic analytic family of maps if
$T\times A$ $\to B$
$(t,a)$ $\mapsto f_{t}(a)$
is a $p$-adic analytic map.
Let $X_{t}\in\mathfrak{X}(A)$ be a vector field for each $t\in T$. We say that $\{X_{t}\}_{t\in T}$ is a $p$-adic analytic family of vector fields if, for all $f\in\Omega_{0}(A)$, $\{X_{t}(f)\}_{t\in T}$ is a $p$-adic analytic family of maps.
Let $\omega_{t}\in\Omega_{k}(A)$ be a $k$-form for each $t\in T$. We say that $\{\omega_{t}\}_{t\in T}$ is a $p$-adic analytic family of $k$-forms if, for every $X_{1},\ldots,X_{k}\in\mathfrak{X}(A)$, $\{\omega_{t}(X_{1},\ldots,X_{k})\}_{t\in T}$ is a $p$-adic analytic family of maps.
Appendix B The real Moser’s Path Method
Moser’s Path Method [29], or Moser’s Trick, is a far-reaching tool used in differential geometry which allows us to find diffeomorphisms between differential forms, provided a vector field satisfying some conditions is found. It can be used to prove Darboux’s Theorem [12], which states that any symplectic form on a $2n$-dimensional manifold is locally symplectomorphic to $\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i},$ where $(x_{1},y_{1},\ldots,x_{n},y_{n})$ are the standard coordinates on $\mathbb{R}^{2n}$ centered at the origin. Darboux’s result can be generalized from neigborhoods of points to neighborhoods of compact submanifolds, this being the content of the Darboux-Weinstein’s Theorem, proved by Alan Weinstein [45, Theorem 4.1]. The precise statement of Moser’s Path Method is as follows.
Let $\ell$ and $k$ be integers such that $0\leqslant k\leqslant\ell$ and $\ell\geqslant 1$. Let $M$ be an $\ell$-dimensional smooth manifold, $m\in M$ and $\{\omega_{t}\}_{t\in[0,1]}$ a family of smooth $k$-forms on $M$. Let $\mathrm{L}$ denote the Lie derivative. Suppose that there exists a smooth time-dependent vector field $\{X_{t}\}_{t\in[0,1]}$ on $M$ such that $\frac{\mathrm{d}}{\mathrm{d}t}\omega_{t}+\mathrm{L}_{X_{t}}\omega_{t}=0$ and $X_{t}(m)=0$ for every $t\in[0,1]$. Then by [29, Theorem 2] there exists an open neighborhood $U$ of $m$ and a smooth family of $p$-adic analytic diffeomorphisms $\{\psi_{t}:U\to\psi_{t}(U)\subset M\}_{t\in[0,1]}$ such that $\psi_{t}^{*}\omega_{t}=\omega_{0}$ for every $t\in[0,1]$.
The condition that $X_{t}(m)=0$ for every $t\in[0,1]$ can be replaced by $M$ being compact, and in this case one obtains a smooth family of diffeomorphisms $\{\psi_{t}:M\to M\}_{t\in[0,1]}$. Compactness is delicate to use in the $p$-adic category and for the purpose of the present paper we work with the first condition ($X_{t}(m)=0$ for every $t\in[0,1]$).
Indeed, Moser’s idea was the following. If either $M$ is compact or $X_{t}(m)=0$ for all $t$, this vector field defines a flow on the entire interval $[0,1]$. Let $\psi_{t}$ be this flow. Then we have that
$\frac{\mathrm{d}}{\mathrm{d}t}\psi_{t}^{*}\omega_{t}=\psi_{t}^{*}\mathrm{L}_{X_{t}}\omega_{t}+\psi_{t}^{*}\frac{\mathrm{d}}{\mathrm{d}t}\omega_{t}=\psi_{t}^{*}\left(\frac{\mathrm{d}}{\mathrm{d}t}\omega_{t}+\mathrm{L}_{X_{t}}\omega_{t}\right)=0.$
This means that $\psi_{t}^{*}\omega_{t}$ is constant, and for $t=0$ is $\omega_{0}$, hence it is always $\omega_{0}$. The trick is very useful in differential geometry and it is used for example in [28, page 121], and is also applicable in more general situations even when $M$ is not compact, see [11, Section 2].
In Theorem A, we extend Moser’s Path Method to the case where the real field is replaced by the field of the $p$-adic numbers $\mathbb{Q}_{p}$. In this new context, Moser’s Path Method cannot be directly applied, because a crucial part of it is the fact that, if the derivative of a smooth function is zero, then the function is constant. In the $p$-adic case this does not hold at all: there are $p$-adic smooth functions from $\mathbb{Q}_{p}$ to $\mathbb{Q}_{p}$, such as $f(x)=\sum_{n=\operatorname{ord}_{p}(x)}^{\infty}a_{n}p^{2n}$ where $x=\sum_{n=\operatorname{ord}_{p}(x)}^{\infty}a_{n}p^{n}$ is the $p$-adic expansion of $x$, with zero derivative everywhere and which are injective.
A first step to address this is to restrict to $p$-adic analytic functions, but even this is not enough: if the derivative of a $p$-adic analytic function is zero everywhere, then the function may be constant only piecewise, not necessarily globally. In order to attain the needed degree of control over our functions, we need them to be “globally” analytic, that is, given by a single power series on an $\ell$-dimensional $p$-adic analytic manifold:
$\sum_{I=(i_{1},\ldots,i_{\ell})\in\mathbb{N}^{\ell}}a_{I}(x_{1}-x_{01})^{i_{1}}\ldots(x_{\ell}-x_{0\ell})^{i_{\ell}}$
which converges in all of its domain. This makes the statement of Moser’s Path Method more technical than its real counterpart. Its proof is also significantly more involved so we divide it into several steps, spread out over sections 2, 3 and 4.
Appendix C Serre’s classification in $p$-adic analytic geometry
In what follows, when we say that two $p$-adic manifolds are “diffeomorphic” we always mean that they are so by means of a $p$-adic analytic diffeomorphism.
Theorem C.1 (Serre [40, Théorème 1], classification of compact $p$-adic analytic manifolds).
Let $d$ be a positive integer. Let $p$ be a prime number. Let $\mathcal{N}_{d,k}$ be the $p$-adic analytic manifold given as follows:
$\mathcal{N}_{d,k}=\bigcup_{i=1}^{k}\mathrm{B}\left(\left(\sum_{j=0}^{c(i)}a_{j}(i)p^{-1-j},0,\ldots,0\right),1\right)$
where
$i=\sum_{j=0}^{c(i)}a_{j}(i)p^{j}$
is the base $p$ expansion of $i$, for $i\in\mathbb{N}$.
  1. (1)
    Let $M$ be a compact $d$-dimensional $p$-adic analytic manifold. Then there exists a positive integer $k$ such that $M$ is diffeomorphic to $\mathcal{N}_{d,k}$.
  2. (2)
    The $p$-adic analytic manifolds $\mathcal{N}_{d,k}$ and $\mathcal{N}_{d,k^{\prime}}$ are diffeomorphic if and only if $k\equiv k^{\prime}\mod p-1$.
In the noncompact case, the classification of Theorem C.1 can be extended to the case where the manifold is second-countable. Indeed, if a $p$-adic analytic manifold is second-countable, then it is paracompact, and by Schneider [38, Proposition 8.7], it is diffeomorphic to a union of balls, which must be a countable union. Hence, any second-countable (in particular, any compact) $d$-dimensional $p$-adic analytic manifold is diffeomorphic either to $(\mathbb{Q}_{p})^{d}$ or to the disjoint union of at most p1p-1 copies of $(\mathbb{Z}_{p})^{d}$.
Theorem C.1 has a consequence in the case where $M$ is endowed with a $p$-adic volume form, that is, a $p$-adic analytic differential $d$-form which never vanishes. The typical cases are subsets of $(\mathbb{Q}_{p})^{d}$, where we have the standard volume form $\mathrm{d}x_{1}\wedge\mathrm{d}x_{2}\wedge\ldots\mathrm{d}x_{d}$, where $(x_{1},\ldots,x_{d})$ are the standard coordinates on $(\mathbb{Q}_{p})^{d}$, and, for $d=2n$, all $2n$-dimensional $p$-adic analytic symplectic manifolds $(M,\omega)$, where $\omega^{n}$ is a volume form. In this case, we can define a $p$-adic volume as the $p$-adic measure with respect to this volume form, see for example [47, p. 14-15] and [35].
Corollary C.2.
Let $M$ and $M^{\prime}$ be compact $d$-dimensional $p$-adic analytic manifolds endowed with a volume form and with the same $p$-adic volume. Then $M$ is diffeomorphic to $M^{\prime}$.